### v2

# Biharmonic nonlinear vector field equations in dimension 44

Abstract
Following the approach of Brezis and Lieb [8], we prove the existence of a ground state solution for the biharmonic nonlinear vector field equations in the limiting case of space dimension $4$. Our results complete those obtained by Mederski and Siemianowski [16] for dimensions $d\geq 5$. We also extend the biharmonic logarithmic Sobolev inequality in [16] to dimension $4$.
keywords:
biharmonic , ground state , elliptic system , logarithmic Sobolev inequality
2020 MSC:
35J91 , 35J20 , 35B38 , 35Qxx
1 Introduction and main results
The study of ground state problems for the Laplacian operator was mainly motivated by the search of stationary states for some nonlinear evolution equations (e.g. of the Klein-Gordon or Schrödinger type), and some other applications in Physics and Biology (e.g. in statistical mechanics [4], cosmology [11], or population dynamics [5, 12]).
Given a potential $G:\mathbb{R}^{m}\to\mathbb{R}$ having a critical point at 0, and satisfying some growth conditions near 0 as well as asymptotically, we consider the semilinear system
$-\Delta u(x)=\nabla G(u(x)),\ u:\mathbb{R}^{d}\to\mathbb{R}^{m},\,d\geq 2,\,m\geq 1,$ (1)
whose associate energy functional is $S(u)=\int_{\mathbb{R}^{d}}\big[\frac{1}{2}|\nabla u|^{2}-G(u)\big]$. The solutions of interest are those converging asymptotically to 0, and minimizing $S$ among all nonzero solutions in an appropriate class. Such a solution is called a ground state. Among the main contributions to solve this problem, we have to mention the seminal paper of Berestycki and Lions [7], which was completed by the work of Berestycki, Gallouët, and Kavian [6], in the two dimensional limiting case, and by the work of Brezis and Lieb [8] in the vector case. The method developed in all these papers is based on a constrained minimization. For more recent developments on these results we refer for instance to the papers [3, 10, 14, 15]. It is also worth noting that in view of the Sobolev embeddings, the case of dimensions $d\geq 3$ differs from the two dimensional limiting case.
Due to the increasing interest for higher order models and their applications (e.g. in nonlinear elasticity [1] and mechanics [17]), it is also relevant to investigate ground state problems for the bilaplacian operator $\Delta^{2}$. This was done by Mederski and Siemienowski [16] for dimensions $d\geq 5$, and as an application they conceived a new biharmonic logarithmic Sobolev inequality (cf. [16, Theorem 1.4]).
The scope of the present paper is to study the limiting case of space dimension $4$. Our approach is based on the method introduced in [8], as well as on some tools borrowed from [16] (e.g. Pohozaev identity in [16, Theorem 1.2]). Following [16], we also extend the biharmonic logarithmic Sobolev inequality in [16, Theorem 1.4] to dimension $4$.
In this context, we consider the system
$\Delta^{2}u(x)=g(u(x)),\ u:\mathbb{R}^{4}\to\mathbb{R}^{m},\,m\geq 1,$ (2)
where the map $g:\mathbb{R}^{m}\to\mathbb{R}^{m}$, $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ [^1] To simplify the notation, we write $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ instead of $g\in L_{loc}^{\infty}(\mathbb{R}^{m};\mathbb{R}^{m})$, and when considering other vector-valued maps belonging to some $L^{p}$ space, we shall in general omit the range., is the gradient (except at $u=0$) of a function $G:\mathbb{R}^{m}\to\mathbb{R}$ satisfying the following assumptions:
$G(0)=0,\ G:\mathbb{R}^{m}\to\mathbb{R}\text{ is continuous, and }G\in C^{1}(\mathbb{R}^{4}\setminus\{0\}),$ (3)
$g(0)=0,\text{ and }g(u)=\nabla G(u),\,\forall u\neq 0,$ (4)
$G(u)<0\ \text{ for }0<|u|\leq\epsilon,\text{ and for some }\epsilon>0,$ (5)
$G(u_{0})>0\ \text{for some}\ u_{0}.$ (6)
That is, the function $G$ has at $u=0$ a local (but not global) maximum equal to 0. We also point out that $G$ need not be differentiable at 0. For instance, we may have G(u)=|u|G(u)=-|u| near 0. The energy functional associated with system (2) is
$S(u)=K(u)-V(u),$ (7)
$\text{where }K(u)=\dfrac{1}{2}\int_{\mathbb{R}^{4}}\left|D^{2}u(x)\right|^{2}dx=\dfrac{1}{2}\sum_{i,j=1}^{4}\int_{\mathbb{R}^{4}}\Big|\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big|^{2}dx,$ (8)
$\text{ and }V(u)=\int_{\mathbb{R}^{4}}G(u(x))dx.$ (9)
Our goal is to show that $S(u)$ admits a minimum in the set of nontrivial solutions $u\neq 0$ to system (2), converging to 0 as $|x|\to\infty$, in some weak sense. Such a least energy solution is called a ground state. Following [8], we shall look for this solution in the class:
$\mathcal{C}=\left\{u\in L^{1}_{loc}(\mathbb{R}^{4})\ \middle|\begin{aligned} &\ \nabla u\in D^{1,2}(\mathbb{R}^{4}),\ G(u)\in L^{1}(\mathbb{R}^{4})\\ &\ \mu(\{|u|>\eta\})<\infty,\ \forall\eta>0\end{aligned}\right\},$
where $\mu$ denotes the Lebesgue measure, and by definition of the space $D^{1,2}$ (cf. [20]), we have
$\nabla u\in D^{1,2}(\mathbb{R}^{4})\Leftrightarrow\nabla u\in L^{4}(\mathbb{R}^{4}),\ D^{2}u\in L^{2}(\mathbb{R}^{4}).$
It is worth noting that $K(u)=\frac{1}{2}\int_{\mathbb{R}^{4}}|\Delta u|^{2}$, provided that $\nabla u\in D^{1,2}(\mathbb{R}^{4})$ (cf. Lemma 11). Thus, $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$. We shall also see in Lemma 1 below, that the class $\mathcal{C}$ is contained in $L^{\infty}(\mathbb{R}^{4})$. Due to this property, and in contrast with the case of the second order system (1), none asymptotic condition on $G$ is required to prove the existence of a ground state solution to (2).
Theorem 1.1.
If assumptions (3)-(6) hold, then
$T=\inf\left\{\int_{\mathbb{R}^{4}}\dfrac{1}{2}|D^{2}u|^{2}:\ u\in\mathcal{C}\ ,\ \int_{\mathbb{R}^{4}}G(u)\geq 0\ ,\ u\neq 0\right\}$ (10)
is achieved by some $\bar{u}\in\mathcal{C}$, with $\bar{u}\neq 0$ and $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, solving (2) in $\mathcal{D}^{\prime}$. Moreover, we have
  • 1.
    $\bar{u}\in C^{3,\alpha}_{loc}\cap W^{4,q}_{loc}$ for any $0<\alpha<1$ and $1\leq q<\infty$,
  • 2.
    $\lim_{|x|\to\infty}\bar{u}(x)=0$,
  • 3.
    and the solution $\bar{u}$ is a ground state, since there holds $0<S(\bar{u})\leq S(v)$ for all $v\in\mathcal{C}$ with $v\neq 0$ which are solutions of (2) in $\mathcal{D}^{\prime}$.
Remark 1.
If $g$ is locally Hölder continuous, then Theorem 1.1 provides a classical solution $\bar{u}\in C^{4}(\mathbb{R}^{4})$.
As an application of Theorem 1.1, we choose the potential
$G(u)=\begin{cases}|u|^{2}\ln|u|&\text{ for }u\neq 0\\ 0&\text{ for }u=0,\end{cases}$
which clearly satisfies assumptions (3)-(6). Moreover, for this choice of $G$, one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$. Following [16], we derive in (12) below the biharmonic logarithmic Sobolev inequality in dimension $4$ (cf. [16, Theorem 1.4] for dimensions $d\geq 5$). It is a generalization for the bilaplacian operator of the classical logarithmic Sobolev inequality
$\ln\left(\dfrac{1}{2\pi e}\int_{\mathbb{R}^{4}}|\nabla u|^{2}\right)\geq\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,\text{ for }u\in H^{1}(\mathbb{R}^{4}),\int_{\mathbb{R}^{4}}|u|^{2}=1,$ (11)
given in [19], which is optimal in the scalar case (cf. [9]).
Theorem 1.2.
Let $\mathcal{C}^{\prime}=\{u\in\mathcal{C}:\int_{\mathbb{R}^{4}}|u|^{2}=1\}$. For any $u\in\mathcal{C}^{\prime}$ there holds
$\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|\Delta u|^{2}\right)=\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)\geq\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,$ (12)
with $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$, where $T>0$ is the minimum in (10). Moreover, the equality in (12) holds provided that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ and $\bar{u}$ is a ground state solution to (2). Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, then there are uniquely determined $\mu>0$ and $r>0$, such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
Remark 2.
Equality case in (12) is achieved for the normalized ground state solutions provided by Theorem 1.1. However, these solutions are not known explicitly, thus we neither know the exact value of the optimal constant in (12). By comparing it with the optimal constant (cf. [9]) of the classical logarithmic Sobolev inequality (11), we can at least obtain that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
The proofs of Theorems 1.1 and 1.2 are given in Sections 2 and 3 respectively, while in the Appendix we gather some technical lemmas.
2 Proof of Theorem 1.1
Let $\{u_{j}\}\subset\mathcal{C}$ be a minimizing sequence for (10), i.e.
$\lim_{j\to\infty}\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u_{j}|^{2}=T.$
Since in the limiting case of dimension $4$, $K(u)$ is invariant under scaling, we may assume that
$\mu(\{|u_{j}|>\varepsilon\})=1.$ (13)
It is clear that the norms $\|D^{2}u_{j}\|_{L^{2}(\mathbb{R}^{4})}$ are uniformly bounded. Therefore, due to the Sobolev imbedding of $D^{1,2}(\mathbb{R}^{4})$ in $L^{4}(\mathbb{R}^{4})$ (cf. [20, Theorem 1.8]), we also obtain that the norms $\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})}$ are uniformly bounded.
Next, we establish some properties of the class $\mathcal{C}$, and the sequence $\{u_{j}\}$.
Lemma 1.
Let $v\in\mathcal{C}$, then $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, $\forall\alpha\in(0,1)$, and it holds $\lim_{|x|\to\infty}v(x)=0$. In addition, the sequence $\{u_{j}\}$ is equibounded as well as equicontinuous in $\mathbb{R}^{4}$, and satisfies
$\int_{\{|u_{j}|>\varepsilon\}}|u_{j}|^{q}\leq C_{q},\ \forall q\in[1,+\infty),\ \forall j\text{ and for some constant }C_{q}.$ (14)
Proof.
We have for every $q\geq 1$ the following estimates.
$\|\psi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\psi\|_{L^{2}(\mathbb{R}^{4})}(\mu(supp\psi))^{1/q},$ (15a)
holding for $\psi\in L^{1}_{loc}(\mathbb{R}^{4})$ such that $\nabla\psi\in L^{2}(\mathbb{R}^{4})$ and $\mu(supp\psi)<\infty$,
$\|\phi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{\frac{1}{q}+\frac{1}{4}},$ (15b)
holding for $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$ and $\mu(supp\phi)<\infty$, where $C^{\prime}_{q}>0$ is a constant. Estimate (15a) can be found in [8, (3.10)], while (15b) follows from (15a) and the Cauchy-Schwarz inequality, since we have $\nabla\phi\in L^{2}(\mathbb{R}^{4})$ and $\|\nabla\phi\|_{L^{2}(\mathbb{R}^{4})}\leq\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{1/4}$.
To see that any $v\in\mathcal{C}$ is bounded, we use the fact that $\nabla v\in L^{4}(\mathbb{R}^{4})$, which implies that $\mu(\{|\nabla v|>1\})<\infty$. Then, applying (15a) to the function $\psi=(|\nabla v|-1)^{+}$, we obtain that
$\gamma_{q}=\int_{|\nabla v|>1}|\nabla v|^{q}$ $\leq\int_{|\nabla v|>1}((|\nabla v|-1)^{+}+1)^{q}$
$\leq(\|\psi\|_{L^{q}(\mathbb{R}^{4})}+\mu(\{|\nabla v|>1\}))^{q}<\infty,$
for any $q\in[1,\infty)$. Consequently, it follows that the integrals
$\int_{B(x,1)}|\nabla v|^{q}$ $=\int_{\{|\nabla v|>1\}\cap B(x,1)}|\nabla v|^{q}+\int_{\{|\nabla v|\leq 1\}\cap B(x,1)}|\nabla v|^{q}$
$\leq\gamma_{q}+\mu(B(0,1)),$
are uniformly bounded on the unit balls of $\mathbb{R}^{4}$. A similar application of (15b) to $\phi=(|v|-1)^{+}$ shows that the integrals $\int_{B(x,1)}|v|^{q}$ are also uniformly bounded. Therefore the $W^{1,q}$ norm of $v$ is uniformly bounded on the unit balls, for any $q\in[1,\infty)$. This implies that $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, for any $\alpha\in(0,1)$.
A consequence of the uniform continuity of $v$ is that the set $\{|v|>\epsilon\}$ is bounded. Indeed, if $|v(x_{n})|>\epsilon$ holds for a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, then there exists $\delta>0$ such that $|v|>\epsilon/2$ holds on the balls $B(x_{n},\delta)$, which by passing to a subsequence can be assumed to be disjoint. Consequently, we obtain that $\mu(\{|v|>\epsilon/2\})=\infty$, a contradiction.
Now, let us assume the existence of a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ satisfying $\lim_{n\to\infty}|x_{n}|=\infty$, and $|v(x_{n})|>\eta$, for some $\eta\in(0,\epsilon)$. According to what precedes, we have $|v(x_{n})|\in[\eta,\epsilon]$ (for $n$ large enough). Then, in view of the uniform continuity of the function $\mathbb{R}^{4}\ni x\mapsto G(v(x))$ and (5), there exists $\delta>0$, such that $|G(v)|>\beta>0$ holds on the balls $B(x_{n},\delta)$, for some $\beta>0$. As previously, the balls $B(x_{n},\delta)$ can be assumed to be disjoint, thus we get $G(v)\notin L^{1}(\mathbb{R}^{4})$, a contradiction. Therefore, we established that $\lim_{|x|\to\infty}v(x)=0$.
Finally, since the measures $\mu(\{|\nabla u_{j}|>1\})$ and $\mu(\{|u_{j}|>\varepsilon\})$ are uniformly bounded, the same arguments applied to the sequences $\psi_{j}=(|\nabla u_{j}|-1)^{+}$ and $\phi_{j}=(|u_{j}|-\epsilon)^{+}$ imply (14), and the fact that the $W^{1,q}$ norms of $\{u_{j}\}$ are uniformly bounded on the unit balls, for any $q\in[1,\infty)$. Thus, the sequence $\{u_{j}\}$ is equibounded and equicontinuous in $\mathbb{R}^{4}$. ∎
At this stage, we recall the following compactness result [13, Lemma 6].
Lemma 2.
Let $v\in L^{4}_{loc}(\mathbb{R}^{4})$, $\nabla v\in L^{4}(\mathbb{R}^{4})$, $\|\nabla v\|_{L^{4}}\leq C$ and $\mu(\{|v|>\varepsilon\})\geq\delta>0$. Then, there exists a shift $T_{y}v(x)=v(x+y)$ such that, for some constant $\beta=\beta(C,\delta,\varepsilon)>0$, $\mu(B\cap\{|T_{y}v|>\varepsilon/2\})>\beta$, where $B=\{x\in\mathbb{R}^{4}:|x|\leq 1\}$.
In view of Lemma 2, we can shift each $u_{j}$ in such a way that $\mu(B\cap\{|T_{x_{j}}u_{j}|>\epsilon/2\})\geq\beta$. Thus, we may assume that $\mu(B\cap\{|u_{j}|>\epsilon/2\})\geq\beta$. On the other hand, the previous estimates imply that we can extract a subsequence still called $\{u_{j}\}$ such that
  1. 1.
    $u_{j}\to u$ uniformly on compact subsets of $\mathbb{R}^{4}$,
  2. 2.
    $\nabla u_{j}\rightharpoonup\nabla u$ in $L^{4}(\mathbb{R}^{4})$,
  3. 3.
    $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$,
  4. 4.
    $\mu(B\cap\{|u|\geq\varepsilon/2\})\geq\beta>0\ \Rightarrow u\not\equiv 0$,
  5. 5.
    $\mu(\{|u|>\varepsilon\})\leq 1$,
  6. 6.
    $u\in\mathcal{C}$.
Properties 4. and 5. of the limit $u$ follow from the convergence in 1., which is established via the theorem of Ascoli and a diagonal argument. To check that $G(u)\in L^{1}$, we use assumption (5) and the inequality $G^{+}(v)\leq k|v|$, holding for $|v|\leq\sup_{j}\|u_{j}\|_{L^{\infty}(\mathbb{R}^{4})}$, and for a constant $k>0$. Then, (14) implies that the integrals $\int_{\mathbb{R}^{4}}G^{+}(u_{j})$ are uniformly bounded, and we deduce from Fatou’s lemma that $\int_{\mathbb{R}^{4}}G^{+}(u)<\infty$. Similarly, since we have $\int_{\mathbb{R}^{4}}G^{-}(u_{j})\leq\int_{\mathbb{R}^{4}}G^{+}(u_{j})$, the integrals $\int_{\mathbb{R}^{4}}G^{-}(u_{j})$ are uniformly bounded, and we obtain as previously that $\int_{\mathbb{R}^{4}}G^{-}(u)<\infty$. Therefore, $G(u)\in L^{1}$. To prove that $u\in\mathcal{C}$, it remains to show that $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$. Indeed, if $\eta\geq\epsilon$, this is clear in view of 5. above. Otherwise, if $0<\eta<\epsilon$, we have
$\mu(\{\eta<|u|\leq\epsilon\})\min\{|G(v)|:\eta\leq|v|\leq\epsilon\}\leq\int_{\eta\leq|u|\leq\epsilon}|G(u)|<\infty,$
thus $\mu(\{\eta<|u|\leq\epsilon\})<\infty\Rightarrow\mu(\{|u|>\eta\})<\infty$.
Following [8], we also define the class
$\mathcal{K}=\left\{\phi:\mathbb{R}^{4}\to\mathbb{R}^{m}|\ \phi\in C_{0}(\mathbb{R}^{4}),\ \nabla\phi\in L^{4},\ D^{2}\phi\in L^{2}\right\},$ (16)
where $C_{0}$ is the space of continuous functions with compact support.
In the sequel, we proceed in several steps as in [8], to construct the solution of Theorem 1.1.
Lemma 3.
Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Then,
$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,$ (17)
where $D^{2}u\cdot D^{2}\phi=\sum_{i,j=1}^{4}\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial^{2}\phi}{\partial x_{i}\partial x_{j}}$.
Proof.
Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Due to the uniform convergence $G(u_{j}+\phi)-G(u_{j})\to G(u+\phi)-G(u)$ on $supp\phi$, we have that
$\int_{\mathbb{R}^{4}}G(u_{j}+\phi)$ $=\int_{supp\phi}G(u_{j}+\phi)+\int_{\mathbb{R}^{4}\setminus supp\phi}G(u_{j})$
$\geq\int_{supp\phi}(G(u_{j}+\phi)-G(u_{j}))>0,$
for $j$ large enough. Thus,
$\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j}+\phi)|^{2}=\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j})|^{2}+\int_{\mathbb{R}^{4}}D^{2}u_{j}\cdot D^{2}\phi+\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq T.$ (18)
Since, $\lim_{j\to\infty}\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j})|^{2}=T$, by letting $j\to\infty$ in (18), we get (17). ∎
Lemma 4.
There exists a constant $C_{G}$ (depending only on $G$) such that if $\phi\in\mathcal{K}$, then
$\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0\ \Rightarrow\ \int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi\geq 0.$ (19)
Proof.
Let $\phi\in\mathcal{K}$. By dominated convergence, one can see that as $t\to 0^{+}$, we have
$\int_{u\neq 0}\big(G(u+t\phi)-G(u)\big)=t\int_{u\neq 0}g(u)\cdot\phi+o(t)=t\int_{\mathbb{R}^{4}}g(u)\cdot\phi+o(t).$ (20)
On the other hand, it is clear that
$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|,\text{ as }t\to 0^{+},$ (21)
holds for a constant $C_{G}>0$ depending only on $G$, since we have the bound $|G(v)|\leq C_{G}|v|$, provided that $|v|\leq\epsilon$.
Now, choose $\phi\in\mathcal{K}$ such that $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0$. Then, for small enough $t>0$, it follows from (20) and (21) that $\int_{\mathbb{R}^{4}}\big(G(u+t\phi)-G(u)\big)>0$. By Lemma 3, we conclude that
$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{t}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,\ \text{ for }0<t\ll 1\Rightarrow\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi\geq 0.$
Lemma 5.
Consider the linear functional $L(\phi)=\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi$. There is some $\phi_{0}\in\mathcal{K}$ such that $L(\phi_{0})\neq 0$ and $\phi_{0}=0$ on the set $\{u=0\}$.
Proof.
Let $\phi_{n}=u\rho_{n}\in\mathcal{K}$, where $\rho_{n}=\rho(x/n)$, and $\rho\in C^{\infty}_{0}(\mathbb{R}^{4})$ is a function such that $0\leq\rho\leq 1$ and
$\rho(x)=\begin{cases}1&\text{ for }|x|\leq 1\\ 0&\text{ for }|x|\geq 2.\end{cases}$ (22)
Let also $A_{n}=\{x\in\mathbb{R}^{4}:n\leq|x|\leq 2n\}$. We assume by contradiction that $L(\phi)=0$, for all $\phi\in\mathcal{K}$ such that $\phi=0$ on $\{u=0\}$. It is clear that $\phi_{n}=0$ on the set $\{u=0\}$. Thus, we have $L(\phi_{n})=L(u\rho_{n})=0$. That is,
$\int_{\mathbb{R}^{4}}\sum_{i,j=1}^{4}\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\Big(\frac{\partial u}{\partial x_{i}}\frac{\partial\rho_{n}}{\partial x_{j}}+\frac{\partial u}{\partial x_{j}}\frac{\partial\rho_{n}}{\partial x_{i}}\Big)+\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big)\\ +\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=0.$
After direct calculations we obtain that
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=\int_{\mathbb{R}^{4}}|D^{2}u|^{2},$
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial u}{\partial x_{i}}\Big)\frac{\partial\rho_{n}}{\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|\nabla u\|_{L^{4}(A_{n})}\to 0\text{, as }n\to\infty,$
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|u\|_{L^{\infty}(\mathbb{R}^{4})}\to 0\text{, as }n\to\infty,$
for a constant $k_{\rho}>0$ depending on $\rho$. This implies that $D^{2}u\equiv 0$, and since $u$ is bounded, we get $u\equiv 0$, which is a contradiction. ∎
Lemmas 3.5 and 3.6 from [8] are applicable to our case without modifications. For convenience of the reader we reproduce in Lemmas 6 and 7 below their proofs.
Lemma 6.
There is a constant $\lambda\geq 0$ such that
$\Big|\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)\Big|\leq C_{G}\int_{u=0}|\phi|,\ \forall\phi\in\mathcal{K}.$ (23)
Proof.
In view of Lemma 5, we can choose $\phi_{0}\in\mathcal{K}$ such that $\phi_{0}=0$ on $\{u=0\}$, and $L(\phi_{0})=-1$. Next, given $\phi\in\mathcal{K}$, we consider the function $\psi=\phi+L(\phi)\phi_{0}+\alpha\phi_{0}$, with $\alpha>0$. It is clear that $L(\psi)=-\alpha<0$, thus by Lemma 4, we have that $\int_{\mathbb{R}^{4}}g(u)\cdot(\phi+L(\phi)\phi_{0}+\alpha\phi_{0})-C_{G}\int_{u=0}|\phi|\leq 0.$ Setting $\lambda=-\int_{\mathbb{R}^{4}}g(u)\cdot\phi_{0}\geq 0$ (by (19)), and letting $\alpha\to 0$, we deduce that $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)-C_{G}\int_{u=0}|\phi|\leq 0$, $\forall\phi\in\mathcal{K}$, from which (23) follows. ∎
Lemma 7.
We have that $\frac{G(u)}{\partial x_{i}}=g(u)\cdot\frac{\partial u}{\partial x_{i}}$ in $\mathcal{D}^{\prime}$.
Proof.
Let $M=\|u\|_{L^{\infty}(\mathbb{R}^{4})}$, and let $G_{k}$ be a smoothing sequence for $G$ so that $G_{k}\to G$ uniformly on $B_{M}=\{v\in\mathbb{R}^{m}:|v|\leq M\}$, and $g_{k}=\nabla G_{k}\to g$ pointwise on $B_{M}\setminus\{0\}$. Moreover, we can ensure that the maps $g_{k}$ are uniformly bounded on $B_{M}$. It is clear that $\frac{\partial G_{k}(u)}{\partial x_{i}}=g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}$. Thus, given $\phi\in C^{\infty}_{0}$ we have $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G_{k}(u)=-\int_{\mathbb{R}^{4}}\phi\big(g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, and passing to the limit we get $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G(u)=-\int_{\mathbb{R}^{4}}\phi\big(g(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, since $\frac{\partial u}{\partial x_{i}}=0$ on the set $\{u=0\}$. ∎
Lemma 8.
Let $\lambda$ be as in Lemma 6. Then, $u$ is a solution of the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $W^{4,q}_{loc}$, $\forall q\in[1,\infty)$, satisfying $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$ and $\int_{\mathbb{R}^{4}}G(u)=0$.
Proof.
The linear functional $M(\phi)=\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)$ restricted to $C^{\infty}_{0}$ satisfies $|M(\phi)|\leq C_{G}\|\phi\|_{L^{1}}$ (cf. (23)), thus by the Riesz Representation theorem, there exists $h\in L^{\infty}(\mathbb{R}^{4})$, $h:\mathbb{R}^{4}\to\mathbb{R}^{m}$, such that $M(\phi)=\int_{\mathbb{R}^{4}}h\cdot\phi$, $\forall\phi\in C^{\infty}_{0}$. Actually, (23) implies that $\int_{\mathbb{R}^{4}}h\cdot\phi=0$, for all $\phi\in C^{\infty}_{0}$ with support in the open set $\{u\neq 0\}$. Consequently, $h$ vanishes on $\{u\neq 0\}$, and denoting by $\chi$ the characteristic function, we have $\lambda\Delta^{2}u+g(u)=h\chi_{u=0}$ in $\mathcal{D}^{\prime}$. Our next claim is that $\lambda>0$. Otherwise, if $\lambda=0$, then $g(u)\equiv 0$. Moreover, in view of Lemma 7, the continuous function $x\mapsto G(u(x))$ is constant. Since $\lim_{|x|\to\infty}G(u(x))=0$, it follows that $G(u)\equiv 0$ and $u\equiv 0$, which is a contradiction. Hence, it holds $\Delta^{2}u=\frac{1}{\lambda}(g(u)-h\chi_{u=0})$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. Finally, since $u\in L^{\infty}$, while $g(u)-h\chi_{u=0}\in L^{\infty}$, we know by Lemma 9 that $u\in W^{4,q}_{loc}$, $\forall q\geq 1$. Therefore, $\Delta^{2}u=0$ a.e. on $\{u=0\}$, and $h=0$ a.e. on $\{u=0\}$. That is, we have $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. In addition, Pohozaev identity (cf. Lemma 10) implies that $\int_{\mathbb{R}^{4}}G(u)=0$, while in view of the weak convergence $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$, we have $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\leq T$. Thus, $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$, and $u$ solves the minimization problem (10). ∎
Setting $\bar{u}(x)=u(\lambda^{-\frac{1}{4}}x)$, we obtain a solution of (2), such that $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, and $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\bar{u}|^{2}=T$ (i.e. $\bar{u}$ still solves the minimization problem (10)). Next, let $v\in\mathcal{C}$ be a solution of (2) in $\mathcal{D}^{\prime}$ such that $v\neq 0$. Then, by Pohozaev identity it holds $\int_{\mathbb{R}^{4}}G(v)=0$, and we have $S(v)=K(v)\geq T=K(\bar{u})=S(\bar{u})$. That is, $\bar{u}$ has all the properties stated in Theorem 1.1. We also point out that if $g$ is locally Hölder continuous, then by elliptic regularity we obtain that $\bar{u}\in C^{4}(\mathbb{R}^{4})$, since the function $x\mapsto g(\bar{u}(x))$ is locally Hölder continuous.
Remark 3.
Any solution $u$ of the minimization problem (10) solves the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, where $\lambda>0$ is the Lagrange multiplier provided by Lemma 6. In addition, there holds $\int_{\mathbb{R}^{4}}G(u)=0$. Indeed, by considering the minimizing sequence $u_{j}=u$, $\forall j$, the arguments of Lemmas 3, 4, 5, 6, 7 and 8 can be reproduced and apply to the minimizer $u$. Thus, the map $\bar{u}(x)=u(rx)$, with $r=\lambda^{-\frac{1}{4}}$, is a ground state solution to (2).
3 Proof of Theorem 1.2
As we mentioned in the Introduction, given the potential $G(u)=|u|^{2}\ln|u|$ (which clearly satisfies assumptions (3)-(6)), one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$. Let $u\in\mathcal{C}^{\prime}$, and set $\mu=e^{-V(u)}$, with $V(u)=\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|$. We compute $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, since $\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}=1$. Consequently, we have the inequality $\frac{\mu^{2}}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\geq T$, which is equivalent to (12). Next, assume that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$, where $\bar{u}$ is a ground state solution to (2). Then, setting $\mu=1/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ one can see that the right hand side of (12) is equal to $\frac{1}{2}\ln(\mu^{2})$, while the left hand side is $V(\mu\bar{u})=\mu^{2}(\ln(\mu)\|\bar{u}\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(\bar{u}))=\ln(\mu)$, since $V(\bar{u})=0$. This is an equality case in (12). Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, i.e. $K(u)=Te^{2V(u)}$ (where $K$ is defined in (7)), then we have $K(\mu u)=T\Leftrightarrow\mu=e^{-V(u)}$. Moreover, for this value of $\mu$ there holds $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, that is, $\mu u$ solves the minimization problem (10). Therefore, in view of Remark 3, there exists a unique $r>0$ such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
To complete the proof of the theorem, it remains to show that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$. To this end, we use (11) (which also holds in the vector case since we have $|\nabla|u||\leq|\nabla u|$, $\forall u\in H^{1}(\mathbb{R}^{4};\mathbb{R}^{m})$), combined with the inequality (cf. [16, Lemma 6.1]):
$\int_{\mathbb{R}^{4}}|\nabla u|^{2}<\Big[\int_{\mathbb{R}^{4}}|\Delta u|^{2}\Big]^{\frac{1}{2}}\Big[\int_{\mathbb{R}^{4}}|u|^{2}\Big]^{\frac{1}{2}},\text{ for }u\in H^{2}(\mathbb{R}^{4}),u\neq 0.$ (24)
In view of Lemma 11, we have $\mathcal{C}^{\prime}\subset H^{2}(\mathbb{R}^{4})$, as well as $\|D^{2}u\|_{L^{2}(\mathbb{R}^{4})}=\|\Delta u\|_{L^{2}(\mathbb{R}^{4})}$, thus we obtain
$\dfrac{1}{2}\ln\left(\dfrac{1}{(2\pi e)^{2}}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)>\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,\text{ for }u\in\mathcal{C}^{\prime},$ (25)
and choosing $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ with $\bar{u}$ a ground state solution to (2), we conclude that
$\dfrac{1}{2}\ln\left(\dfrac{1}{(2\pi e)^{2}}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)>\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|=\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right).$ (26)
This proves that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
4 Some lemmas applicable to the solutions of (2)
Here we gather some lemmas applicable to the solutions of (2). We first recall a regularity result for the nonhomogeneous biharmonic equation that can be derived from the $L^{p}$ estimates in [18, (2.6)].
Lemma 9.
Let $1<p<\infty$, and $\Omega\subset\mathbb{R}^{n}$ an open set. If $u\in L^{p}_{loc}(\Omega)$ and $\Delta^{2}u\in L^{p}_{loc}(\Omega)$, then $u\in W^{4,p}_{loc}(\Omega)$.
Next, we state Pohozaev identity for solutions to (2) belonging to $\mathcal{C}$. This identity was established in [16, Theorem 1.2] for dimensions $d\geq 5$. Its proof is identical in dimension $4$.
Lemma 10.
Let $v\in\mathcal{C}$ be a solution to (2) in $\mathcal{D}^{\prime}$. Then, $\int_{\mathbb{R}^{4}}G(v)=0$.
Proof.
We notice that since $v$ as well as $g(v)$ are bounded in $\mathbb{R}^{4}$, it follows from Lemma 9 that $v\in W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q\geq 1$. Then, we reproduce the arguments in the proof of [16, Theorem 1.2]. ∎
Finally, we establish that $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$.
Lemma 11.
Let $u\in L^{1}_{loc}(\mathbb{R}^{4})$ be such that $\nabla u\in L^{4}(\mathbb{R}^{4})$, and $D^{2}u\in L^{2}(\mathbb{R}^{4})$. Then, we have $\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=\int_{\mathbb{R}^{4}}|\Delta u|^{2}$. If in addition $u\in L^{2}(\mathbb{R}^{4})$, then it holds $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla u|^{2}=-\int_{\mathbb{R}^{4}}u\cdot\Delta u$.
Proof.
(i) Let $\phi\in C^{\infty}(\mathbb{R}^{4})$ be such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $D^{2}\phi\in L^{2}(\mathbb{R}^{4})$. By integrating by parts we obtain
$\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}=\lim_{r\to\infty}\sum_{i,j=1}^{4}\int_{B_{r}}\partial^{2}_{ij}\phi\cdot\partial^{2}_{ij}\phi$
$=$ $\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[-\int_{B_{r}}\partial_{i}\phi\cdot\partial^{3}_{ijj}\phi+\int_{\partial B_{r}}\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi\Big]$
$=$ $\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[\int_{B_{r}}\partial^{2}_{ii}\phi\cdot\partial^{2}_{jj}\phi+\int_{\partial B_{r}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big]$
$=$ $\int_{\mathbb{R}^{4}}|\Delta\phi|^{2}+\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[\int_{\partial B_{r}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big],$
where $B_{r}=\{x\in\mathbb{R}^{4}:|x|<1\}$, and $\nu=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})$ is the outward unit normal vector at $\partial B_{r}$. Moreover, since $|\nabla\phi||D^{2}\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}=0$. Therefore, we conclude that
$\Big|\int_{\partial B_{r_{k}}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
as $k\to\infty$, and $\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}=\int_{\mathbb{R}^{4}}|\Delta\phi|^{2}$. Now, if $u\in L^{1}_{loc}(\mathbb{R}^{4})$ is such that $\nabla u\in L^{4}(\mathbb{R}^{4})$, and $D^{2}u\in L^{2}(\mathbb{R}^{4})$, we consider a sequence $\rho_{n}$ of mollifiers. Setting $\phi_{n}=u\star\rho_{n}\in C^{\infty}(\mathbb{R}^{4})$, we have $\partial_{i}\phi_{n}=\partial_{i}u\star\rho_{n}\to\partial_{i}u$ in $L^{4}(\mathbb{R}^{4})$, as well as $\partial^{2}_{ij}\phi_{n}=\partial^{2}_{ij}u\star\rho_{n}\to\partial^{2}_{ij}u$ in $L^{2}(\mathbb{R}^{4})$. Consequently, it holds $\int_{\mathbb{R}^{4}}|D^{2}\phi_{n}|^{2}=\int_{\mathbb{R}^{4}}|\Delta\phi_{n}|^{2}$, and as $n\to\infty$, we get $\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=\int_{\mathbb{R}^{4}}|\Delta u|^{2}$.
(ii) To prove the second statement of the lemma, we consider a function $\phi\in C^{\infty}(\mathbb{R}^{4})\cap L^{2}(\mathbb{R}^{4})$ such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $D^{2}\phi\in L^{2}(\mathbb{R}^{4})$. As previously, we integrate by parts and obtain
$\int_{B_{r}}|\nabla\phi|^{2}=-\int_{B_{r}}\phi\cdot\Delta\phi+\int_{\partial B_{r}}\frac{\partial\phi}{\partial\nu}\cdot\phi.$
On the one hand, it is clear that $\lim_{r\to\infty}\int_{B_{r}}\phi\cdot\Delta\phi=\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. On the other hand, we notice that since $|\nabla\phi||\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi=0$. Therefore, we have
$\Big|\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
as $k\to\infty$, and $\lim_{k\to\infty}\int_{B_{r_{k}}}|\nabla\phi|^{2}=-\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. This proves that $\nabla\phi\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla\phi|^{2}=-\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. Finally, if $u\in L^{2}(\mathbb{R}^{4})$ is such that $\nabla u\in L^{4}(\mathbb{R}^{4})$, and $D^{2}u\in L^{2}(\mathbb{R}^{4})$, we proceed as previously by mollification. Setting $\phi_{n}=u\star\rho_{n}\in C^{\infty}(\mathbb{R}^{4})$, we have that $\phi_{n}\to u$ as well as $\Delta\phi_{n}=\Delta u\star\rho_{n}\to\Delta u$ in $L^{2}(\mathbb{R}^{4})$, while $\partial_{i}\phi_{n}=\partial_{i}u\star\rho_{n}\to\partial_{i}u$ in $L^{4}(\mathbb{R}^{4})$ and a.e. in $\mathbb{R}^{4}$ (up to subsequence). Consequently, it holds $\int_{\mathbb{R}^{4}}|\nabla\phi_{n}|^{2}=-\int_{\mathbb{R}^{4}}\phi_{n}\cdot\Delta\phi_{n}$, and as $n\to\infty$, we get that $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla u|^{2}=-\int_{\mathbb{R}^{4}}u\cdot\Delta u$.
5 Conclusion
The method of constrained minimization applies to prove the existence of a ground state solution for both systems (1) and (2), in space dimensions greater or equal than that of the limiting case (i.e. $d\geq 2$ for (1), and $d\geq 4$ for (2)). In contrast with the second order system (1), none asymptotic condition on the potential $G$ is required for the fourth order system (2), since for the latter, the relevant class $\cal C$ of functions is contained in $L^{\infty}$. We also point out that if the space dimension is smaller than that of the limiting case, then the problem of finding nontrivial solutions converging asympototically to 0 is different. In space dimension $d=1$, such solutions of (1) are the homoclinic orbits (cf. for instance [2]). On the other hand, the construction of such solutions of (2) in space dimensions $d=1,2,3$, is to a large extent an open problem.
Acknowledgements
This research project is implemented within the framework of H.F.R.I call “Basic research Financing (Horizontal support of all Sciences)” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union - NextGenera tionEU (H.F.R.I. Project Number: 016097).

### v3

@@
 in nonlinear elasticity [4] and mechanics [26]), it is also relevant to investigate ground state problems for the bilaplacian operator $\Delta^{2}$.
-This was done by Mederski and Siemienowski [16] for dimensions $d\geq 5$, and as an application they conceived a new biharmonic logarithmic Sobolev inequality (cf.
-[16, Theorem 1.4]).
+This was done by Mederski and Siemienowski [23] for dimensions $d\geq 5$, and as an application they derived a biharmonic logarithmic Sobolev inequality (cf.
+[23, Theorem 1.4]).
 The scope of the present paper is to study the limiting case of space dimension $4$.
 Our approach is based on the method introduced in [9], as well as on some tools borrowed from [23] (e.g.
-Pohozaev identity in [16, Theorem 1.2]).
-Following [16], we also extend the biharmonic logarithmic Sobolev inequality in [16, Theorem 1.4] to dimension $4$.
+Pohozaev identity in [23, Theorem 1.2]).
+Following [23], we also extend the biharmonic logarithmic Sobolev inequality in [23, Theorem 1.4] to dimension $4$.
 In this context, we consider the system
 $\Delta^{2}u(x)=g(u(x)),\ u:\mathbb{R}^{4}\to\mathbb{R}^{m},\,m\geq 1,$ (2)
-where the map $g:\mathbb{R}^{m}\to\mathbb{R}^{m}$, $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ [^1] To simplify the notation, we write $g\in L^{\infty}_{loc}(\mathbb{R}^{m})$ instead of $g\in L_{loc}^{\infty}(\mathbb{R}^{m};\mathbb{R}^{m})$, and when considering other vector-valued maps belonging to some $L^{p}$ space, we shall in general omit the range., is the gradient (except at $u=0$) of a function $G:\mathbb{R}^{m}\to\mathbb{R}$ satisfying the following assumptions:
+where the map $g:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is the gradient (except at $u=0$) of a function $G:\mathbb{R}^{m}\to\mathbb{R}$ satisfying the following assumptions:
 $G(0)=0,\ G:\mathbb{R}^{m}\to\mathbb{R}\text{ is continuous, and }G\in C^{1}(\mathbb{R}^{4}\setminus\{0\}),$ (3)
@@
 $G(u)<0\ \text{ for }0<|u|\leq\epsilon,\text{ and for some }\epsilon>0,$ (5)
-$G(u_{0})>0\ \text{for some}\ u_{0}.$ (6)
+$G(u_{0})>0\ \text{for some}\ u_{0},$ (6)
+$\text{$\forall b>0,\ \forall\delta>0,\ \exists C_{\delta,b}>0$ such that }|g(u)|\leq C_{\delta,b}+\delta e^{b|u|^{4/3}},\quad\forall u\in\mathbb{R}^{4}.$ (7)
 That is, the function $G$ has at $u=0$ a local (but not global) maximum equal to 0.
@@
 For instance, we may have G(u)=|u|G(u)=-|u| near 0.
+On the other hand, it follows from (7) that
+$\lim_{|u|\to\infty}\dfrac{|g(u)|}{e^{a|u|^{4/3}}}=0\ \text{and}\ \lim_{|u|\to\infty}\dfrac{|G(u)|}{e^{a|u|^{4/3}}}=0,\ \forall a>0.$ (8)
+We shall see later that this growth condition is required to obtain compactness properties from the embedding of $W^{1,4}(\mathbb{R}^{4})$ in the Orlicz space $L_{A}(\mathbb{R}^{4})$, where the function $A$ is defined in (24).
 The energy functional associated with system (2) is
 $S(u)=K(u)-V(u),$ (9)
-$\text{where }K(u)=\dfrac{1}{2}\int_{\mathbb{R}^{4}}\left|D^{2}u(x)\right|^{2}dx=\dfrac{1}{2}\sum_{i,j=1}^{4}\int_{\mathbb{R}^{4}}\Big|\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big|^{2}dx,$ (8)
-$\text{ and }V(u)=\int_{\mathbb{R}^{4}}G(u(x))dx.$ (9)
+$\text{where }K(u)=\dfrac{1}{2}\int\left|D^{2}u(x)\right|^{2}dx=\dfrac{1}{2}\sum_{i,j=1}^{4}\int\Big|\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big|^{2}dx,$ (10)
+$\text{ and }V(u)=\int G(u(x))dx.$ (11)
 Our goal is to show that $S(u)$ admits a minimum in the set of nontrivial solutions $u\neq 0$ to system (2), converging to 0 as $|x|\to\infty$, in some weak sense.
@@
 Thus, $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$.
-We shall also see in Lemma 1 below, that the class $\mathcal{C}$ is contained in $L^{\infty}(\mathbb{R}^{4})$.
-Due to this property, and in contrast with the case of the second order system (1), none asymptotic condition on $G$ is required to prove the existence of a ground state solution to (2).
 Theorem 1.1.
@@
   • 2.
-    $\lim_{|x|\to\infty}\bar{u}(x)=0$,
+    $\lim_{|x|\to\infty}\bar{u}(x)=\lim_{|x|\to\infty}|\nabla\bar{u}(x)|=0$,
   • 3.
@@
 Remark 1.
-If $g$ is locally Hölder continuous, then Theorem 1.1 provides a classical solution $\bar{u}\in C^{4}(\mathbb{R}^{4})$.
+If $g$ is locally Hölder continuous, then Theorem 1.1 provides a classical solution $\bar{u}\in C^{4}(\mathbb{R}^{4})$, since the function $x\mapsto g(\bar{u}(x))$ is locally Hölder continuous.
 As an application of Theorem 1.1, we choose the potential
@@
 which clearly satisfies assumptions (3)-(7).
-Moreover, for this choice of $G$, one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$.
+Moreover, for this choice of $G$, one can see that $\mathcal{C}=\{u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m}):\,G^{-}(u)\in L^{1}(\mathbb{R}^{4})\}$ (cf.
+Remark 6).
 Following [23], we derive in (14) below the biharmonic logarithmic Sobolev inequality in dimension $4$ (cf.
-[16, Theorem 1.4] for dimensions $d\geq 5$).
+[23, Theorem 1.4] for dimensions $d\geq 5$).
 It is a generalization for the bilaplacian operator of the classical logarithmic Sobolev inequality
@@
 Theorem 1.2.
-Let $\mathcal{C}^{\prime}=\{u\in\mathcal{C}:\int_{\mathbb{R}^{4}}|u|^{2}=1\}$. For any $u\in\mathcal{C}^{\prime}$ there holds
+For any $u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m})$ such that $\int_{\mathbb{R}^{4}}|u|^{2}=1$, there holds
 $\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|\Delta u|^{2}\right)=\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)\geq\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,$ (14)
-with $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$, where $T>0$ is the minimum in (10).
+with $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$, where $T>0$ is the minimum in (12).
 Moreover, the equality in (14) holds provided that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ and $\bar{u}$ is a ground state solution to (2).
-Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, then there are uniquely determined $\mu>0$ and $r>0$, such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
+Conversely, if the equality in (14) holds for $u$, then there are uniquely determined $\mu>0$ and $r>0$, such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
+Remark 2.
+If $u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m})$, we have $(|u|^{2}\ln|u|)^{+}\in L^{1}(\mathbb{R}^{4})$ (cf.
+Remark 6), and hence the right hand side of (14) is well defined, although it could be -$\infty$ if $\int_{\mathbb{R}^{4}}(|u|^{2}\ln|u|)^{-}=+\infty$.
 Remark 3.
@@
 By comparing it with the optimal constant (cf.
-[9]) of the classical logarithmic Sobolev inequality (11), we can at least obtain that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
-The proofs of Theorems 1.1 and 1.2 are given in Sections 2 and 3 respectively, while in the Appendix we gather some technical lemmas.
+[10]) of the classical logarithmic Sobolev inequality (13), we can at least obtain that $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$.
+Logarithmic Sobolev type inequalities appear in several equivalent formulations.
+The seminal inequality by Gross [18] was established in the Gaussian setting.
+Weissler’s approach [27] gives the Lebesgue analogue with sharp constant which under the unit normalization $\|u\|_{L^{2}(\mathbb{R}^{4})}=1$ leads to the form (13) (see also [20]).
+In a series of papers [12, 13], Cotsiolis and Tavoularis proved sharp Sobolev type inequalities with explicit best constants for higher order fractional derivatives in the range $d>2s$.
+Most relevant here is their higher order fractional derivative logarithmic result [14], which for every $s>0$ and every $\alpha>0$ gives a sharp inequality with explicit optimal constants: for $u\in H^{s}(\mathbb{R}^{d})$,
+$\int_{\mathbb{R}^{d}}|u|^{2}\ln\!\Big(\frac{|u|^{2}}{\|u\|_{2}^{2}}\Big)\,dx+\Big(d+\frac{d}{s}\ln\alpha+\ln\!\Big(s\,\frac{\Gamma(d/2)}{\Gamma(d/(2s))}\Big)\Big)\|u\|_{2}^{2}\\ \;\leq\;\frac{\alpha^{2}}{\pi^{s}}\|(-\Delta)^{s/2}u\|_{2}^{2}.$ (15)
+In particular, for $s=1$ one recovers the Gross inequality (13) by setting $\|u\|_{2}=1$ and minimizing in $\alpha$.
+Finally, Mederski and Siemianowski (for $d\geq 5$) derived a scalar biharmonic logarithmic Sobolev inequality tied to the bilaplacian, i.e. for $u\in D^{2,2}(\mathbb{R}^{d})$ with $\|u\|_{2}=1$,
+$\int_{\mathbb{R}^{d}}|u|^{2}\ln|u|\,dx\;\leq\;\frac{d}{8}\ln\!\Big(C\int_{\mathbb{R}^{d}}|\Delta u|^{2}\,dx\Big),$ (16)
+with some $0<C<\big(\frac{2}{\pi ed}\big)^{2}$.
+Moreover, they relate the equality to normalized ground state solutions of the associated biharmonic problem and note that the exact optimal constant is not explicitly computable.
+Note that, the case $s=2$ in (15) corresponds to the logarithmic inequality controlled by $\|\Delta u\|_{2}^{2}$, matching the inequality above.
+The proofs of Theorems 1.1 and 1.2 are given in Sections 2 and 3 respectively, while in Section 4 we gather some technical lemmas.
 2 Proof of Theorem 1.1
@@
 Since in the limiting case of dimension $4$, $K(u)$ is invariant under scaling, we may assume that
-$\mu(\{|u_{j}|>\varepsilon\})=1.$ (13)
+[…]
 It is clear that the norms $\|D^{2}u_{j}\|_{L^{2}(\mathbb{R}^{4})}$ are uniformly bounded.
 Therefore, due to the Sobolev imbedding of $D^{1,2}(\mathbb{R}^{4})$ in $L^{4}(\mathbb{R}^{4})$ (cf.
-[20, Theorem 1.8]), we also obtain that the norms $\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})}$ are uniformly bounded.
-Next, we establish some properties of the class $\mathcal{C}$, and the sequence $\{u_{j}\}$.
+[28, Thm.
+1.8]), we also obtain that the norms $\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})}$ are uniformly bounded.
+Next, we show that $\{u_{j}\}\subset L^{q}_{loc}(\mathbb{R}^{4})$, $\forall q\in[1,+\infty)$, and estimate the $L^{q}$ norm of $u_{j}$.
 Lemma 1.
-Let $v\in\mathcal{C}$, then $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, $\forall\alpha\in(0,1)$, and it holds $\lim_{|x|\to\infty}v(x)=0$.
-In addition, the sequence $\{u_{j}\}$ is equibounded as well as equicontinuous in $\mathbb{R}^{4}$, and satisfies
-$\int_{\{|u_{j}|>\varepsilon\}}|u_{j}|^{q}\leq C_{q},\ \forall q\in[1,+\infty),\ \forall j\text{ and for some constant }C_{q}.$ (14)
+We have that
+$\int_{|u_{j}|>\varepsilon/2}|u_{j}|^{q}\leq C_{q},\ \forall q\in[1,+\infty),\ \forall j\text{ and for some constant }C_{q}.$ (18)
 Proof.
-We have for every $q\geq 1$ the following estimates.
-$\|\psi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\psi\|_{L^{2}(\mathbb{R}^{4})}(\mu(supp\psi))^{1/q},$ (15a)
-holding for $\psi\in L^{1}_{loc}(\mathbb{R}^{4})$ such that $\nabla\psi\in L^{2}(\mathbb{R}^{4})$ and $\mu(supp\psi)<\infty$,
-$\|\phi\|_{L^{q}(\mathbb{R}^{4})}\leq C^{\prime}_{q}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{\frac{1}{q}+\frac{1}{4}},$ (15b)
-holding for $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$ and $\mu(supp\phi)<\infty$, where $C^{\prime}_{q}>0$ is a constant.
-Estimate (15a) can be found in [8, (3.10)], while (15b) follows from (15a) and the Cauchy-Schwarz inequality, since we have $\nabla\phi\in L^{2}(\mathbb{R}^{4})$ and $\|\nabla\phi\|_{L^{2}(\mathbb{R}^{4})}\leq\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\phi))^{1/4}$.
-To see that any $v\in\mathcal{C}$ is bounded, we use the fact that $\nabla v\in L^{4}(\mathbb{R}^{4})$, which implies that $\mu(\{|\nabla v|>1\})<\infty$.
-Then, applying (15a) to the function $\psi=(|\nabla v|-1)^{+}$, we obtain that
-$\gamma_{q}=\int_{|\nabla v|>1}|\nabla v|^{q}$ $\leq\int_{|\nabla v|>1}((|\nabla v|-1)^{+}+1)^{q}$
-$\leq(\|\psi\|_{L^{q}(\mathbb{R}^{4})}+\mu(\{|\nabla v|>1\}))^{q}<\infty,$
-for any $q\in[1,\infty)$. Consequently, it follows that the integrals
-$\int_{B(x,1)}|\nabla v|^{q}$ $=\int_{\{|\nabla v|>1\}\cap B(x,1)}|\nabla v|^{q}+\int_{\{|\nabla v|\leq 1\}\cap B(x,1)}|\nabla v|^{q}$
-$\leq\gamma_{q}+\mu(B(0,1)),$
-are uniformly bounded on the unit balls of $\mathbb{R}^{4}$.
-A similar application of (15b) to $\phi=(|v|-1)^{+}$ shows that the integrals $\int_{B(x,1)}|v|^{q}$ are also uniformly bounded.
-Therefore the $W^{1,q}$ norm of $v$ is uniformly bounded on the unit balls, for any $q\in[1,\infty)$.
-This implies that $v\in L^{\infty}(\mathbb{R}^{4})\cap C^{\alpha}(\mathbb{R}^{4})$, for any $\alpha\in(0,1)$.
-A consequence of the uniform continuity of $v$ is that the set $\{|v|>\epsilon\}$ is bounded.
-Indeed, if $|v(x_{n})|>\epsilon$ holds for a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, then there exists $\delta>0$ such that $|v|>\epsilon/2$ holds on the balls $B(x_{n},\delta)$, which by passing to a subsequence can be assumed to be disjoint.
-Consequently, we obtain that $\mu(\{|v|>\epsilon/2\})=\infty$, a contradiction.
-Now, let us assume the existence of a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ satisfying $\lim_{n\to\infty}|x_{n}|=\infty$, and $|v(x_{n})|>\eta$, for some $\eta\in(0,\epsilon)$.
-According to what precedes, we have $|v(x_{n})|\in[\eta,\epsilon]$ (for $n$ large enough).
-Then, in view of the uniform continuity of the function $\mathbb{R}^{4}\ni x\mapsto G(v(x))$ and (5), there exists $\delta>0$, such that $|G(v)|>\beta>0$ holds on the balls $B(x_{n},\delta)$, for some $\beta>0$.
-As previously, the balls $B(x_{n},\delta)$ can be assumed to be disjoint, thus we get $G(v)\notin L^{1}(\mathbb{R}^{4})$, a contradiction.
-Therefore, we established that $\lim_{|x|\to\infty}v(x)=0$.
-Finally, since the measures $\mu(\{|\nabla u_{j}|>1\})$ and $\mu(\{|u_{j}|>\varepsilon\})$ are uniformly bounded, the same arguments applied to the sequences $\psi_{j}=(|\nabla u_{j}|-1)^{+}$ and $\phi_{j}=(|u_{j}|-\epsilon)^{+}$ imply (14), and the fact that the $W^{1,q}$ norms of $\{u_{j}\}$ are uniformly bounded on the unit balls, for any $q\in[1,\infty)$.
-Thus, the sequence $\{u_{j}\}$ is equibounded and equicontinuous in $\mathbb{R}^{4}$.
-∎
-At this stage, we recall the following compactness result [13, Lemma 6].
+Our first claim is that given $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$ and $\mu(supp\,\phi)<\infty$, we have
+$\|\phi\|_{L^{q}(\mathbb{R}^{4})}$ $\leq k_{q}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}(\mu(supp\,\phi))^{\frac{1}{q}},\ \forall q\geq\frac{4}{3},$ (19)
+and for a constant $k_{q}>0$. To establish (19), we follow the steps of the proof of [9, inequality (3.10)]. Indeed, it holds (cf. [25, Theorem p. 125]):
+$\|\phi\|_{L^{\frac{4}{3}}(\mathbb{R}^{4})}$ $\leq k\|\nabla\phi\|_{L^{1}(\mathbb{R}^{4})},\ \forall\phi\in C^{1}_{0}(\mathbb{R}^{4}),$ (20)
+for a constant $k>0$. Next, by applying (20) to the functions $\phi^{n}$, $n=2,3,\ldots$, we get by induction
+$\|\phi\|_{L^{\frac{4n}{3}}(\mathbb{R}^{4})}$ $\leq k_{n}\|\nabla\phi\|_{L^{1}(\mathbb{R}^{4})}^{\frac{1}{n}}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}^{\frac{n-1}{n}},$ (21)
+for a constant $k_{n}>0$, and by interpolation
+$\|\phi\|_{L^{q}(\mathbb{R}^{4})}$ $\leq k_{q}\|\nabla\phi\|_{L^{1}(\mathbb{R}^{4})}^{\frac{4}{3q}}\|\nabla\phi\|_{L^{4}(\mathbb{R}^{4})}^{1-\frac{4}{3q}},\ \forall q\geq\frac{4}{3},$ (22)
+for a constant $k_{q}>0$.
+Moreover, smoothing by convolution, we notice that (22) still holds for functions $\phi\in L^{\infty}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $\phi$ has compact support.
+On the other hand, a standard cut-off and truncation argument shows the validity of (22) for functions $\phi\in L^{1}_{loc}(\mathbb{R}^{4})$, such that $\nabla\phi\in L^{4}(\mathbb{R}^{4})$, and $\mu(supp\,\phi)<\infty$.
+Finally, (19) follows from (22) by Hölder’s inequality.
+To conclude the proof of the lemma, we apply (19) to $\phi=(|u_{j}|-\varepsilon/2)^{+}$, and obtain
+$\|(|u_{j}|-\varepsilon/2)^{+}\|_{L^{q}(\mathbb{R}^{4})}\leq k_{q}\|\nabla u_{j}\|_{L^{4}(\mathbb{R}^{4})},$
+which implies (18). ∎
+At this stage, we recall the following compactness result [19, Lemma 6].
 Lemma 2.
-Let $v\in L^{4}_{loc}(\mathbb{R}^{4})$, $\nabla v\in L^{4}(\mathbb{R}^{4})$, $\|\nabla v\|_{L^{4}}\leq C$ and $\mu(\{|v|>\varepsilon\})\geq\delta>0$.
-Then, there exists a shift $T_{y}v(x)=v(x+y)$ such that, for some constant $\beta=\beta(C,\delta,\varepsilon)>0$, $\mu(B\cap\{|T_{y}v|>\varepsilon/2\})>\beta$, where $B=\{x\in\mathbb{R}^{4}:|x|\leq 1\}$.
-In view of Lemma 2, we can shift each $u_{j}$ in such a way that $\mu(B\cap\{|T_{x_{j}}u_{j}|>\epsilon/2\})\geq\beta$.
-Thus, we may assume that $\mu(B\cap\{|u_{j}|>\epsilon/2\})\geq\beta$.
-On the other hand, the previous estimates imply that we can extract a subsequence still called $\{u_{j}\}$ such that
-  1. 1.
-    $u_{j}\to u$ uniformly on compact subsets of $\mathbb{R}^{4}$,
-  2. 2.
+Let $v\in L^{4}_{loc}(\mathbb{R}^{4})$, $\nabla v\in L^{4}(\mathbb{R}^{4})$, $\|\nabla v\|_{L^{4}}\leq C$ and $\mu(\{|v|>\varepsilon/2\})\geq\delta>0$.
+Then, there exists a shift $T_{y}v(x)=v(x+y)$ such that, for some constant $\beta=\beta(C,\delta,\varepsilon)>0$, $\mu(B\cap\{|T_{y}v|>\varepsilon/4\})>\beta$, where $B=\{x\in\mathbb{R}^{4}:|x|\leq 1\}$.
+In view of Lemma 2, we can shift each $u_{j}$ in such a way that $\mu(B\cap\{|T_{x_{j}}u_{j}|>\epsilon/4\})\geq\beta$.
+Thus, we may assume that $\mu(B\cap\{|u_{j}|>\epsilon/4\})\geq\beta$.
+On the other hand, Lemma 1 implies that given a bounded set $Q\subset\mathbb{R}^{4}$, the norms $\|u_{j}\|_{L^{q}(Q)}$ are uniformly bounded.
+Consequently, we can extract a subsequence still called $\{u_{j}\}$ such that
+  1. (i)
+    $u_{j}\to u$ in $L^{q}(Q)$, $\forall q\in[1,+\infty)$, $\forall Q\subset\mathbb{R}^{4}$ bounded,
+  2. (ii)
+    $u_{j}\to u$ a.e. in $\mathbb{R}^{4}$,
+  3. (iii)
     $\nabla u_{j}\rightharpoonup\nabla u$ in $L^{4}(\mathbb{R}^{4})$,
-  3. 3.
+  4. (iv)
     $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$,
-  4. 4.
-    $\mu(B\cap\{|u|\geq\varepsilon/2\})\geq\beta>0\ \Rightarrow u\not\equiv 0$,
-  5. 5.
-    $\mu(\{|u|>\varepsilon\})\leq 1$,
-  6. 6.
-    $u\in\mathcal{C}$.
-Properties 4.
-and 5.
-of the limit $u$ follow from the convergence in 1., which is established via the theorem of Ascoli and a diagonal argument.
-To check that $G(u)\in L^{1}$, we use assumption (5) and the inequality $G^{+}(v)\leq k|v|$, holding for $|v|\leq\sup_{j}\|u_{j}\|_{L^{\infty}(\mathbb{R}^{4})}$, and for a constant $k>0$.
-Then, (14) implies that the integrals $\int_{\mathbb{R}^{4}}G^{+}(u_{j})$ are uniformly bounded, and we deduce from Fatou’s lemma that $\int_{\mathbb{R}^{4}}G^{+}(u)<\infty$.
+  5. (v)
+    $\mu(B\cap\{|u|\geq\varepsilon/4\})\geq\beta>0\ \Rightarrow u\not\equiv 0$,
+  6. (vi)
+    $\mu(\{|u|>\varepsilon/2\})\leq 1$,
+  7. (vii)
+    $u\in\mathcal{C}$,
+Properties (v) and (vi) of the limit $u$ follow from the a.e. convergence in (ii). To check that $G(u)\in L^{1}$, we use assumption (8) and the embedding
+$W^{1,4}(\mathbb{R}^{4})\hookrightarrow L_{A}(\mathbb{R}^{4}),$ (23)
+where $L_{A}$ is the Orlicz space corresponding to the function
+$A(t)=\exp(t^{4/3})-1-t^{4/3}-\dfrac{1}{2}t^{8/3}$ (24)
+(cf.
+Trudinger’s theorem in [1, §8.29]).
+Since the sequence $\{(|u_{j}|-\epsilon/2)^{+}\}$ is uniformly bounded in $W^{1,4}(\mathbb{R}^{4})$, it follows that there exists $k_{1}>0$, such that the integrals $\int_{\mathbb{R}^{4}}A(k_{1}(|u_{j}|-\epsilon/2)^{+})$ are uniformly bounded.
+Moreover, it holds $G^{+}(u_{j})\leq k_{2}A(k_{1}(|u_{j}|-\epsilon/2)^{+})$, for a constant $k_{2}>0$ (cf.
+(5) and (8)).
+Consequently, the integrals $\int_{\mathbb{R}^{4}}G^{+}(u_{j})$ are uniformly bounded, and we deduce from Fatou’s lemma that $\int_{\mathbb{R}^{4}}G^{+}(u)<\infty$.
 Similarly, since we have $\int_{\mathbb{R}^{4}}G^{-}(u_{j})\leq\int_{\mathbb{R}^{4}}G^{+}(u_{j})$, the integrals $\int_{\mathbb{R}^{4}}G^{-}(u_{j})$ are uniformly bounded, and we obtain as previously that $\int_{\mathbb{R}^{4}}G^{-}(u)<\infty$.
 Therefore, $G(u)\in L^{1}$.
-To prove that $u\in\mathcal{C}$, it remains to show that $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$.
-Indeed, if $\eta\geq\epsilon$, this is clear in view of 5.
-above.
-Otherwise, if $0<\eta<\epsilon$, we have
-$\mu(\{\eta<|u|\leq\epsilon\})\min\{|G(v)|:\eta\leq|v|\leq\epsilon\}\leq\int_{\eta\leq|u|\leq\epsilon}|G(u)|<\infty,$
-thus $\mu(\{\eta<|u|\leq\epsilon\})<\infty\Rightarrow\mu(\{|u|>\eta\})<\infty$.
+To prove that $u\in\mathcal{C}$, it remains to show that $\mu(\{|u|>\alpha\})<\infty$, $\forall\alpha>0$.
+Indeed, if $\alpha\geq\epsilon/2$, this is clear in view of (vi) above.
+Otherwise, if $0<\alpha<\epsilon/2$, we have in view of hypothesis (5):
+$\mu(\{\alpha<|u|\leq\epsilon/2\})\min\{|G(v)|:\alpha\leq|v|\leq\epsilon/2\}\leq\int_{\alpha\leq|u|\leq\epsilon/2}|G(u)|<\infty,$
+thus $\mu(\{\alpha<|u|\leq\epsilon/2\})<\infty\Rightarrow\mu(\{|u|>\alpha\})<\infty$. This completes the proof of property (vii) above.
 Following [9], we also define the class
-$\mathcal{K}=\left\{\phi:\mathbb{R}^{4}\to\mathbb{R}^{m}|\ \phi\in C_{0}(\mathbb{R}^{4}),\ \nabla\phi\in L^{4},\ D^{2}\phi\in L^{2}\right\},$ (16)
-where $C_{0}$ is the space of continuous functions with compact support.
+$\mathcal{K}:=\left\{\phi:\mathbb{R}^{4}\to\mathbb{R}^{m}\ |\ \phi\in L^{1}_{loc},\ \nabla\phi\in L^{4},\ D^{2}\phi\in L^{2},\ \mu(supp\,\phi)<\infty\right\}.$ (25)
+Remark 4.
+Due to (19) we have that $\mathcal{K}\subset L^{q}(\mathbb{R}^{4})$, $\forall q\in[1,\infty)$.
 In the sequel, we proceed in several steps as in [9], to construct the solution of Theorem 1.1.
 Lemma 3.
-Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Then,
-$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,$ (17)
-where $D^{2}u\cdot D^{2}\phi=\sum_{i,j=1}^{4}\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial^{2}\phi}{\partial x_{i}\partial x_{j}}$.
+Let $\phi\in\mathcal{K}$ be such that
+$\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0.$ (26)
+Then,
+$\int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi+\dfrac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq 0,\text{ where }D^{2}u\cdot D^{2}\phi=\sum_{i,j=1}^{4}\partial^{2}_{ij}u\cdot\partial^{2}_{ij}\phi.$ (27)
+Note.
+We shall see in the proof of Lemma 3 that $G(u+\phi)\in L^{1}(\mathbb{R}^{4})$, $\forall\phi\in\mathcal{K}$.
 Proof.
-Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Due to the uniform convergence $G(u_{j}+\phi)-G(u_{j})\to G(u+\phi)-G(u)$ on $supp\phi$, we have that
-$\int_{\mathbb{R}^{4}}G(u_{j}+\phi)$ $=\int_{supp\phi}G(u_{j}+\phi)+\int_{\mathbb{R}^{4}\setminus supp\phi}G(u_{j})$
-$\geq\int_{supp\phi}(G(u_{j}+\phi)-G(u_{j}))>0,$
-for $j$ large enough. Thus,
+Let $\phi\in\mathcal{K}$. We first notice that the sequence $\{(|u_{j}+\phi|-\epsilon/2)^{+}\}$ is uniformly bounded in $L^{4}(\mathbb{R}^{4})$. Indeed, this follows from the inequality
+$\int_{|u_{j}+\phi|>\epsilon/2}|u_{j}+\phi|^{4}$ $\leq\int_{|u_{j}|>\epsilon/2}|u_{j}+\phi|^{4}+\int_{supp\,\phi\cap\{|u_{j}|\leq\epsilon/2\}}|u_{j}+\phi|^{4},$
+combined with (18), and the inclusion $\mathcal{K}\subset L^{4}(\mathbb{R}^{4})$.
+Moreover, one can see that the sequence $\{(|u_{j}+\phi|-\epsilon/2)^{+}\}$ is also uniformly bounded in $W^{1,4}(\mathbb{R}^{4})$.
+Therefore, there exists $k_{1}>0$, such that the integrals $\int_{\mathbb{R}^{4}}A(2k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})$ are uniformly bounded.
+Next, we check that the sequence $\{G(u_{j}+\phi)\}$ is uniformly integrable on $\mathbb{R}^{4}$.
+To prove this, we use the bounds $|G(u)|\leq k_{2}+A(k_{1}(|u|-\epsilon/2)^{+})$, and $|A(|u|)|^{2}\leq k_{3}+A(2|u|)$, holding in $\mathbb{R}^{m}$ for some constants $k_{2},k_{3}>0$.
+Then, given $\omega\subset\mathbb{R}^{4}$, with $\mu(\omega)<\infty$, we obtain
+$\int_{\omega}|G(u_{j}+\phi)|$ $\leq k_{2}\mu(\omega)+\int_{\omega}A(k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})$
+$\leq k_{2}\mu(\omega)+(\mu(\omega))^{\frac{1}{2}}\Big(\int_{\omega}|A(k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})|^{2}\Big)^{\frac{1}{2}}$
+$\leq k_{2}\mu(\omega)+(\mu(\omega))^{\frac{1}{2}}\Big(k_{3}\mu(\omega)+\int_{\mathbb{R}^{4}}A(2k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})\Big)^{\frac{1}{2}},$
+which implies the uniform integrability of $\{G(u_{j}+\phi)\}$ on $\mathbb{R}^{4}$.
+Consequently, since $G(u_{j}+\phi)\to G(u+\phi)$ a.e.
+in $\mathbb{R}^{4}$, we deduce that $G(u+\phi)$ is integrable on $\omega=supp\,\phi$, as well as on $\mathbb{R}^{4}$.
+In addition, we have $\lim_{j\to\infty}\int_{supp\,\phi}G(u_{j}+\phi)=\int_{supp\,\phi}G(u+\phi)$.
+The same argument also proves that the sequence $\{G(u_{j})\}$ is uniformly integrable on $\mathbb{R}^{4}$, and that $\lim_{j\to\infty}\int_{supp\,\phi}G(u_{j})=\int_{supp\,\phi}G(u)$.
+Therefore, we have $\lim_{j\to\infty}\int_{\mathbb{R}^{4}}(G(u_{j}+\phi)-G(u_{j}))=\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))$.
+Finally, assuming (26), $u_{j}+\phi\in\mathcal{C}$ satisfies $\int_{\mathbb{R}^{4}}(G(u_{j}+\phi)-G(u_{j}))>0$, as well as $\int_{\mathbb{R}^{4}}G(u_{j}+\phi)>0$, for $j$ large enough.
+Thus, it follows that
 $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j}+\phi)|^{2}=\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}(u_{j})|^{2}+\int_{\mathbb{R}^{4}}D^{2}u_{j}\cdot D^{2}\phi+\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}\geq T.$ (28)
@@
 $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0\ \Rightarrow\ \int_{\mathbb{R}^{4}}D^{2}u\cdot D^{2}\phi\geq 0.$ (29)
+Note.
+We shall see in the proof of Lemma 4 that $g(u)\cdot\phi\in L^{1}(\mathbb{R}^{4})$, $\forall\phi\in\mathcal{K}$.
 Proof.
-Let $\phi\in\mathcal{K}$. By dominated convergence, one can see that as $t\to 0^{+}$, we have
+Let $\phi\in\mathcal{K}$. By dominated convergence, one can see that $g(u)\cdot\phi\in L^{1}(\mathbb{R}^{4})$, and as $t\to 0^{+}$, we have
 $\int_{u\neq 0}\big(G(u+t\phi)-G(u)\big)=t\int_{u\neq 0}g(u)\cdot\phi+o(t)=t\int_{\mathbb{R}^{4}}g(u)\cdot\phi+o(t).$ (30)
-On the other hand, it is clear that
-$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|,\text{ as }t\to 0^{+},$ (21)
-holds for a constant $C_{G}>0$ depending only on $G$, since we have the bound $|G(v)|\leq C_{G}|v|$, provided that $|v|\leq\epsilon$.
-Now, choose $\phi\in\mathcal{K}$ such that $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0$.
-Then, for small enough $t>0$, it follows from (20) and (21) that $\int_{\mathbb{R}^{4}}\big(G(u+t\phi)-G(u)\big)>0$.
+Indeed, it is clear that $\lim_{t\to 0}\frac{G(u(x)+t\phi(x))-G(u(x))}{t}=g(u(x))\cdot\phi(x)$, provided that $u(x)\neq 0$.
+On the other hand, since $(|u|-\epsilon/2)^{+}+|\phi|\in W^{1,4}(\mathbb{R}^{4})$ (cf.
+(19)), there exists $k_{1}>0$, such that $\int_{\mathbb{R}^{4}}A(2k_{1}[(|u|-\epsilon/2)^{+}+|\phi|])<\infty$.
+Next, in view of the bound $|g(u)|\leq k_{2}+A(k_{1}(|u|-\epsilon/2)^{+})$ and the monotonicity of $A$, we deduce that
+$\forall s,t\in[0,1]:\ |g(u+st\phi)|$ $\leq k_{2}+A(k_{1}(|u+st\phi|-\epsilon/2)^{+})$
+$\leq k_{2}+A(k_{1}[(|u|-\epsilon/2)^{+}+|\phi|]).$
+Consequently, we obtain the inequality
+$\Big|\frac{G(u(x)+t\phi(x))-G(u(x))}{t}\Big|=\Big|\int_{0}^{1}\big(g(u(x)+st\phi(x))\cdot\phi(x)\big)ds\Big|\\ \leq k_{2}|\phi(x)|+A(k_{1}[(|u(x)|-\epsilon/2)^{+}+|\phi(x)|])|\phi(x)|,$
+whose right hand side is an integrable function (since $A(k_{1}[(|u(x)|-\epsilon/2)^{+}+|\phi(x)|])\in L^{2}(supp\,\phi)$ in view of the bound $|A(|u|)|^{2}\leq k_{3}+A(2|u|)$).
+This proves that $g(u)\cdot\phi\in L^{1}(\mathbb{R}^{4})$ and (30) holds.
+Our next claim is that
+$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|+o(t),\text{ as }t\to 0^{+}.$ (31)
+To justify this estimate, we notice that given $\phi\in\mathcal{K}\subset W^{1,4}(\mathbb{R}^{4})$, there exists $k_{4}>0$ such that $A(k_{4}|\phi|)\in L^{1}(\mathbb{R}^{4})$.
+In view of the bound $|G(u)|\leq C_{G}|u|+k_{5}A(k_{4}|u|)$, holding in $\mathbb{R}^{m}$ for a constant $C_{G}>0$ depending only on $G$, and a constant $k_{5}>0$ depending also on $k_{4}$ (cf.
+(7)), it follows that
+$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|+k_{5}\int_{u=0}A(k_{4}t|\phi|).$
+Finally, using the fact that $A(k_{4}t|\phi|)\leq t^{4}A(k_{4}|\phi|)$ holds for $0<t<1$, we deduce (31).
+Now, choose $\phi\in\mathcal{K}$ such that
+$\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0.$
+Then, for small enough $t>0$, it follows from (30) and (31) that
+$\int_{\mathbb{R}^{4}}\big(G(u+t\phi)-G(u)\big)>0.$
 By Lemma 3, we conclude that
@@
 Proof.
-Let $\phi_{n}=u\rho_{n}\in\mathcal{K}$, where $\rho_{n}=\rho(x/n)$, and $\rho\in C^{\infty}_{0}(\mathbb{R}^{4})$ is a function such that $0\leq\rho\leq 1$ and
+Let $\phi_{n}=\rho_{n}u\in\mathcal{K}$, where $\rho_{n}=\rho(x/n)$, and $\rho\in C^{\infty}_{0}(\mathbb{R}^{4})$ is a function such that $0\leq\rho\leq 1$ and
 $\rho(x)=\begin{cases}1&\text{ for }|x|\leq 1\\ 0&\text{ for }|x|\geq 2.\end{cases}$ (32)
@@
 It is clear that $\phi_{n}=0$ on the set $\{u=0\}$.
-Thus, we have $L(\phi_{n})=L(u\rho_{n})=0$.
+Thus, we have $L(\phi_{n})=L(\rho_{n}u)=0$.
 That is,
-$\int_{\mathbb{R}^{4}}\sum_{i,j=1}^{4}\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\Big(\frac{\partial u}{\partial x_{i}}\frac{\partial\rho_{n}}{\partial x_{j}}+\frac{\partial u}{\partial x_{j}}\frac{\partial\rho_{n}}{\partial x_{i}}\Big)+\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big)\\ +\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=0.$
+$\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}+\sum_{i,j=1}^{4}\big(\partial^{2}_{ij}u\cdot\big(\partial_{j}\rho_{n}\,\partial_{i}u+\partial_{i}\rho_{n}\,\partial_{j}u\big)+\big(u\cdot\partial^{2}_{ij}u\big)\partial^{2}_{ij}\rho_{n}\big)=0.$
 After direct calculations we obtain that
 $\lim_{n\to\infty}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=\int_{\mathbb{R}^{4}}|D^{2}u|^{2},$
-$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial u}{\partial x_{i}}\Big)\frac{\partial\rho_{n}}{\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|\nabla u\|_{L^{4}(A_{n})}\to 0\text{, as }n\to\infty,$
-$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|u\|_{L^{\infty}(\mathbb{R}^{4})}\to 0\text{, as }n\to\infty,$
-for a constant $k_{\rho}>0$ depending on $\rho$. This implies that $D^{2}u\equiv 0$, and since $u$ is bounded, we get $u\equiv 0$, which is a contradiction. ∎
-Lemmas 3.5 and 3.6 from [8] are applicable to our case without modifications. For convenience of the reader we reproduce in Lemmas 6 and 7 below their proofs.
+$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\big|\big(\partial^{2}_{ij}u\cdot\partial_{i}u\big)\partial_{j}\rho_{n}\big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|\nabla u\|_{L^{4}(A_{n})}\to 0\text{, as }n\to\infty,$
+$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\big|\big(u\cdot\partial^{2}_{ij}u\big)\partial^{2}_{ij}\rho_{n}\big|$ $\leq\frac{k_{\rho}}{n}\|D^{2}u\|_{L^{2}(A_{n})}\|u\|_{L^{4}(A_{n})}$
+$\leq\frac{k_{\rho}}{n}\|D^{2}u\|_{L^{2}(A_{n})}\big(\|(|u|-\frac{\epsilon}{2})^{+}\|_{L^{4}(A_{n})}+\frac{\epsilon}{2}\mu(A_{n})^{\frac{1}{4}}\big)$
+[…]
+where $k_{\rho}>0$ is a constant depending on $\rho$.
+This implies that $D^{2}u\equiv 0$ i.e.
+$u$ is affine, and since $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$, we get $u\equiv 0$, which is a contradiction.
+∎
+Lemmas 3.5 from [9] is applicable to our case without modifications. For convenience of the reader we reproduce in Lemmas 6 below its proof.
 Lemma 6.
@@
 Lemma 7.
-We have that $\frac{G(u)}{\partial x_{i}}=g(u)\cdot\frac{\partial u}{\partial x_{i}}$ in $\mathcal{D}^{\prime}$.
+We have that
+$\frac{\partial[G(u)]}{\partial x_{i}}=g(u)\cdot\frac{\partial u}{\partial x_{i}}\text{ in }\mathcal{D}^{\prime}.$ (34)
+Note.
+It is clear that $G(u)\in L^{1}(\mathbb{R}^{4})$.
+On the other hand, we shall see in the proof of Lemma 7 that $g(u)\cdot\frac{\partial u}{\partial x_{i}}\in L^{1}_{loc}(\mathbb{R}^{4})$, thus (34) makes sense.
 Proof.
-Let $M=\|u\|_{L^{\infty}(\mathbb{R}^{4})}$, and let $G_{k}$ be a smoothing sequence for $G$ so that $G_{k}\to G$ uniformly on $B_{M}=\{v\in\mathbb{R}^{m}:|v|\leq M\}$, and $g_{k}=\nabla G_{k}\to g$ pointwise on $B_{M}\setminus\{0\}$.
-Moreover, we can ensure that the maps $g_{k}$ are uniformly bounded on $B_{M}$.
-It is clear that $\frac{\partial G_{k}(u)}{\partial x_{i}}=g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}$.
-Thus, given $\phi\in C^{\infty}_{0}$ we have $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G_{k}(u)=-\int_{\mathbb{R}^{4}}\phi\big(g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, and passing to the limit we get $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G(u)=-\int_{\mathbb{R}^{4}}\phi\big(g(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, since $\frac{\partial u}{\partial x_{i}}=0$ on the set $\{u=0\}$.
-∎
+As in [9, Lemma 3.6], we choose a smoothing sequence $\{G_{j}\}\subset C^{\infty}(\mathbb{R}^{m})$ for $G$, satisfying the following properties:
+  • (i)
+    $G_{j}\to G$, as $j\to\infty$, pointwise on $\mathbb{R}^{m}$,
+  • (ii)
+    $g_{j}=\nabla G_{j}\to g$, as $j\to\infty$, pointwise on $\mathbb{R}^{m}\setminus\{0\}$,
+  • (iii)
+    $\forall k_{2}>0$, $\exists k_{3}>0$ such that the inequalities $|g_{j}(u)|\leq k_{3}+e^{k_{2}|u|^{\frac{4}{3}}}$, and $|G_{j}(u)|\leq k_{3}+e^{k_{2}|u|^{\frac{4}{3}}}$, hold on $\mathbb{R}^{m}$ for all $j$.
+Next, given an arbitrary open ball $B\subset\mathbb{R}^{4}$, we can find a sequence $\{v_{n}\}\subset C^{\infty}(\mathbb{R}^{4})$, such that $v_{n}$ converges to $u$ in $W^{1,4}(B)$, while $v_{n}$ and $\nabla v_{n}$ converge almost everywhere respectively to $u$ and $\nabla u$.
+In view of the Trudinger inequality [1, §8.27], there exists a constant $k_{1}>0$ such that the functions $e^{k_{1}|v_{n}|^{\frac{4}{3}}}$ are uniformly bounded in $L^{1}(B)$.
+Choosing $k_{2}=\frac{k_{1}}{2}$, it follows from (iii) above, that the functions $g_{j}(v_{n})$ and $G_{j}(v_{n})$ are uniformly bounded in $L^{2}(B)$.
+Consequently, given a test function $\phi\in C^{\infty}_{0}(B)$ and $i\in\{1,\ldots,4\}$, one can see that the sequences $\{G_{j}(v_{n})\partial_{i}\phi\}_{n}$ and $\{(g_{j}(v_{n})\cdot\partial_{i}v_{n})\phi\}_{n}$ are uniformly integrable over $B$.
+Since there holds
+$\int_{B}G_{j}(v_{n})\partial_{i}\phi=-\int_{B}(g_{j}(v_{n})\cdot\partial_{i}v_{n})\phi,$
+by letting $n\to\infty$, we deduce form the Vitali convergence theorem that
+$\int_{B}G_{j}(u)\partial_{i}\phi=-\int_{B}(g_{j}(u)\cdot\partial_{i}u)\phi.$
+Finally, by letting $j\to\infty$, and using the uniform integrability over $B$ of the sequences $\{G_{j}(u)\partial_{i}\phi\}_{j}$ and $\{(g_{j}(u)\cdot\partial_{i}u)\phi\}_{j}$, as well as the fact that $\partial_{i}u=0$ a.e.
+on the set $\{u=0\}$, we conclude again from the Vitali convergence theorem that
+$\int_{B}G(u)\partial_{i}\phi=-\int_{B}(g(u)\cdot\partial_{i}u)\phi.$
 Lemma 8.
 Let $\lambda$ be as in Lemma 6.
-Then, $u$ is a solution of the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $W^{4,q}_{loc}$, $\forall q\in[1,\infty)$, satisfying $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$ and $\int_{\mathbb{R}^{4}}G(u)=0$.
+Then, $\lambda>0$ and $u$ is a distributional solution of the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$, belonging to $W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q\in[1,\infty)$.
 Proof.
@@
 (33)), thus by the Riesz Representation theorem, there exists $h\in L^{\infty}(\mathbb{R}^{4})$, $h:\mathbb{R}^{4}\to\mathbb{R}^{m}$, such that $M(\phi)=\int_{\mathbb{R}^{4}}h\cdot\phi$, $\forall\phi\in C^{\infty}_{0}$.
-Actually, (23) implies that $\int_{\mathbb{R}^{4}}h\cdot\phi=0$, for all $\phi\in C^{\infty}_{0}$ with support in the open set $\{u\neq 0\}$.
+Furthermore, (33) implies that $|\int_{\mathbb{R}^{4}}h\cdot\phi|\leq C_{G}\int_{u=0}|\phi|$, for all $\phi\in C^{\infty}_{0}$, and thus for all $\phi\in L^{1}(\mathbb{R}^{4})$.
 Consequently, $h$ vanishes on $\{u\neq 0\}$, and denoting by $\chi$ the characteristic function, we have $\lambda\Delta^{2}u+g(u)=h\chi_{u=0}$ in $\mathcal{D}^{\prime}$.
 Our next claim is that $\lambda>0$.
-Otherwise, if $\lambda=0$, then $g(u)\equiv 0$.
-Moreover, in view of Lemma 7, the continuous function $x\mapsto G(u(x))$ is constant.
-Since $\lim_{|x|\to\infty}G(u(x))=0$, it follows that $G(u)\equiv 0$ and $u\equiv 0$, which is a contradiction.
+Indeed, if we assume that $\lambda=0$, then we obtain successively $g(u)\equiv 0$, and $\nabla(G(u))\equiv 0$ (by Lemma 7), from which it follows that $G(u)\equiv 0$, since $G(u)\in L^{1}(\mathbb{R}^{4})$.
+In particular we would have either $u(x)=0$ or $|u(x)|\geq\epsilon$, for a.e.
+$x\in\mathbb{R}^{4}$ (cf.
+(5)), and $\nabla(\min(|u|,\epsilon)\equiv 0$ in $\mathbb{R}^{4}$, which implies that the function $\min(|u|,\epsilon)$ is constant.
+This is excluded, since $u\in\mathcal{C}$ and $u\neq 0$.
 Hence, it holds $\Delta^{2}u=\frac{1}{\lambda}(g(u)-h\chi_{u=0})$ in $\mathcal{D}^{\prime}$, with $\lambda>0$.
-Finally, since $u\in L^{\infty}$, while $g(u)-h\chi_{u=0}\in L^{\infty}$, we know by Lemma 9 that $u\in W^{4,q}_{loc}$, $\forall q\geq 1$.
-Therefore, $\Delta^{2}u=0$ a.e.
+Finally, we check that given $q>1$ and a bounded set $Q\subset\mathbb{R}^{4}$, we have $g(u)\in L^{q}(Q)$.
+To see this, we use the fact that $\int_{\mathbb{R}^{4}}A(k_{1}(|u|-\epsilon/2)^{+})<\infty$, for a constant $k_{1}>0$ (since $(|u|-\epsilon/2)^{+}\in W^{1,4}(\mathbb{R}^{4})$), together with the bounds $|g(u)|\leq k_{2}+A(\frac{k_{1}}{q}(|u|-\epsilon/2)^{+})$ and $|A(|u|)|^{q}\leq k_{3}+A(q|u|)$ holding in $\mathbb{R}^{m}$ for constants $k_{2},k_{3}>0$.
+Therefore, we have $u\in L^{q}_{loc}(\mathbb{R}^{4})$ as well as $g(u)-h\chi_{u=0}\in L^{q}_{loc}(\mathbb{R}^{4})$, $\forall q>1$, and in view of Lemma 9 we deduce that $u\in W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q>1$.
+This implies that $\Delta^{2}u=0$ a.e.
 on $\{u=0\}$, and $h=0$ a.e.
@@
 That is, we have $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, with $\lambda>0$.
-In addition, Pohozaev identity (cf.
+∎
+Finally, Pohozaev identity (cf.
 Lemma 10) implies that $\int_{\mathbb{R}^{4}}G(u)=0$, while in view of the weak convergence $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$, we have $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\leq T$.
 Thus, $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$, and $u$ solves the minimization problem (12).
-∎
-Setting $\bar{u}(x)=u(\lambda^{-\frac{1}{4}}x)$, we obtain a solution of (2), such that $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, and $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\bar{u}|^{2}=T$ (i.e.
+By setting $\bar{u}(x)=u(\lambda^{-\frac{1}{4}}x)$, we obtain a solution of (2), such that $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, and $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\bar{u}|^{2}=T$ (i.e.
 $\bar{u}$ still solves the minimization problem (12)).
 Next, let $v\in\mathcal{C}$ be a solution of (2) in $\mathcal{D}^{\prime}$ such that $v\neq 0$.
-Then, by Pohozaev identity it holds $\int_{\mathbb{R}^{4}}G(v)=0$, and we have $S(v)=K(v)\geq T=K(\bar{u})=S(\bar{u})$.
-That is, $\bar{u}$ has all the properties stated in Theorem 1.1.
-We also point out that if $g$ is locally Hölder continuous, then by elliptic regularity we obtain that $\bar{u}\in C^{4}(\mathbb{R}^{4})$, since the function $x\mapsto g(\bar{u}(x))$ is locally Hölder continuous.
+Then, it holds $\int_{\mathbb{R}^{4}}G(v)=0$ in view of Pohozaev identity, and we have $S(v)=K(v)\geq T=K(\bar{u})=S(\bar{u})$.
+That is, $\bar{u}$ is a ground state solution.
+We also point out that the asymptotic convergence of $\bar{u}$ and $\nabla\bar{u}$ to 0 follows from Lemma 10.
 Remark 5.
@@
 3 Proof of Theorem 1.2
-As we mentioned in the Introduction, given the potential $G(u)=|u|^{2}\ln|u|$ (which clearly satisfies assumptions (3)-(6)), one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$.
-Let $u\in\mathcal{C}^{\prime}$, and set $\mu=e^{-V(u)}$, with $V(u)=\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|$.
+Given the potential $G(u)=|u|^{2}\ln|u|$ (which clearly satisfies assumptions (3)-(7)), we know by Remark 6 that $\mathcal{C}=\{u\in H^{2}(\mathbb{R}^{4}):\,G^{-}(u)\in L^{1}(\mathbb{R}^{4})\}$.
+Let $\mathcal{C}^{\prime}=\{u\in\mathcal{C}:\int_{\mathbb{R}^{4}}|u|^{2}=1\}$, let $u\in\mathcal{C}^{\prime}$, and set $\mu=e^{-V(u)}$, with $V(u)=\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|$.
 We compute $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, since $\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}=1$.
@@
 Next, assume that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$, where $\bar{u}$ is a ground state solution to (2).
-Then, setting $\mu=1/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ one can see that the right hand side of (12) is equal to $\frac{1}{2}\ln(\mu^{2})$, while the left hand side is $V(\mu\bar{u})=\mu^{2}(\ln(\mu)\|\bar{u}\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(\bar{u}))=\ln(\mu)$, since $V(\bar{u})=0$.
+Then, setting $\mu=1/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ one can see that the left hand side of (14) is equal to $\frac{1}{2}\ln(\mu^{2})$, while the right hand side is $V(\mu\bar{u})=\mu^{2}(\ln(\mu)\|\bar{u}\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(\bar{u}))=\ln(\mu)$, since $V(\bar{u})=0$.
 This is an equality case in (14).
@@
 Therefore, in view of Remark 5, there exists a unique $r>0$ such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
-To complete the proof of the theorem, it remains to show that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
+To complete the proof of the theorem, it remains to show that $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$.
 To this end, we use (13) (which also holds in the vector case since we have $|\nabla|u||\leq|\nabla u|$, $\forall u\in H^{1}(\mathbb{R}^{4};\mathbb{R}^{m})$), combined with the inequality (cf.
-[16, Lemma 6.1]):
-$\int_{\mathbb{R}^{4}}|\nabla u|^{2}<\Big[\int_{\mathbb{R}^{4}}|\Delta u|^{2}\Big]^{\frac{1}{2}}\Big[\int_{\mathbb{R}^{4}}|u|^{2}\Big]^{\frac{1}{2}},\text{ for }u\in H^{2}(\mathbb{R}^{4}),u\neq 0.$ (24)
-In view of Lemma 11, we have $\mathcal{C}^{\prime}\subset H^{2}(\mathbb{R}^{4})$, as well as $\|D^{2}u\|_{L^{2}(\mathbb{R}^{4})}=\|\Delta u\|_{L^{2}(\mathbb{R}^{4})}$, thus we obtain
+[23, Lemma 6.1]):
+$\int_{\mathbb{R}^{4}}|\nabla u|^{2}<\Big[\int_{\mathbb{R}^{4}}|\Delta u|^{2}\Big]^{\frac{1}{2}}\Big[\int_{\mathbb{R}^{4}}|u|^{2}\Big]^{\frac{1}{2}},\text{ for }u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m}),u\neq 0.$ (35)
+In view of Lemma 11, we have $\|D^{2}u\|_{L^{2}(\mathbb{R}^{4})}=\|\Delta u\|_{L^{2}(\mathbb{R}^{4})}$, thus we obtain
 $\dfrac{1}{2}\ln\left(\dfrac{1}{(2\pi e)^{2}}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)>\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|,\text{ for }u\in\mathcal{C}^{\prime},$ (36)
@@
 $\dfrac{1}{2}\ln\left(\dfrac{1}{(2\pi e)^{2}}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right)>\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|=\dfrac{1}{2}\ln\left(\dfrac{1}{2T}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\right).$ (37)
-This proves that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.
+This proves that $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$.
 4 Some lemmas applicable to the solutions of (2)
-Here we gather some lemmas applicable to the solutions of (2).
-We first recall a regularity result for the nonhomogeneous biharmonic equation that can be derived from the $L^{p}$ estimates in [18, (2.6)].
+Here, we gather some lemmas applicable to the solutions of (2).
+We first recall a regularity result for the nonhomogeneous biharmonic equation that can be derived from the $L^{p}$ estimates in [24, (2.6)].
 Lemma 9.
 Let $1<p<\infty$, and $\Omega\subset\mathbb{R}^{n}$ an open set. If $u\in L^{p}_{loc}(\Omega)$ and $\Delta^{2}u\in L^{p}_{loc}(\Omega)$, then $u\in W^{4,p}_{loc}(\Omega)$.
-Next, we state Pohozaev identity for solutions to (2) belonging to $\mathcal{C}$. This identity was established in [16, Theorem 1.2] for dimensions $d\geq 5$. Its proof is identical in dimension $4$.
+Next, we examine the regularity of solutions of (2) belonging to $\mathcal{C}$, and state the corresponding Pohozaev identity.
+This identity was established in [23, Theorem 1.2] for dimensions $d\geq 5$.
+Its proof is identical in dimension $4$.
 Lemma 10.
-Let $v\in\mathcal{C}$ be a solution to (2) in $\mathcal{D}^{\prime}$. Then, $\int_{\mathbb{R}^{4}}G(v)=0$.
+Assuming that assumptions (3)-(7) hold, let $v\in\mathcal{C}$ be a ditributional solution of (2). Then, we have
+  • (i)
+    $v\in C^{3,\alpha}_{loc}\cap W^{4,q}_{loc}$ for any $0<\alpha<1$ and $1\leq q<\infty$,
+  • (ii)
+    $\lim_{|x|\to\infty}v(x)=\lim_{|x|\to\infty}|\nabla v(x)|=0$,
+  • (iii)
+    $\int_{\mathbb{R}^{4}}G(v)=0$ (Pohozaev identity).
 Proof.
-We notice that since $v$ as well as $g(v)$ are bounded in $\mathbb{R}^{4}$, it follows from Lemma 9 that $v\in W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q\geq 1$.
-Then, we reproduce the arguments in the proof of [16, Theorem 1.2].
-∎
+(i) Let $v\in\mathcal{C}$ be a ditributional solution of (2).
+Since, $\mu(\{|v|>\epsilon/2\})<\infty$, we can see as in the proof of Lemma 1, that inequality (18) applies to $v$, for every $q>1$.
+In particular, given $q>1$, the norms $\|v\|_{L^{q}(B(x,1))}$ are uniformly bounded on unit balls ($\forall x\in\mathbb{R}^{4}$).
+Similarly, the norms $\|g(v)\|_{L^{q}(B(x,1))}$ are uniformly bounded on unit balls.
+To see this, we use the fact that $\int_{\mathbb{R}^{4}}A(k_{1}(|v|-\epsilon/2)^{+})<\infty$, for a constant $k_{1}>0$ (since $(|v|-\epsilon/2)^{+}\in W^{1,4}(\mathbb{R}^{4})$), together with the bounds $|g(v)|\leq k_{2}+A(\frac{k_{1}}{q}(|v|-\epsilon/2)^{+})$ and $|A(v)|^{q}\leq k_{3}+A(qv)$ holding in $\mathbb{R}^{m}$ for constants $k_{2},k_{3}>0$.
+Therefore, it follows from Lemma 9 that $v\in W^{4,q}_{loc}$, $\forall q>1$, and from the Sobolev embeddings that $v\in C^{3,\alpha}_{loc}$, $\forall\alpha\in(0,1)$.
+(ii) Let $w=\Delta v\in L^{2}(\mathbb{R}^{4})$.
+Another consequence of what precedes is that the norms $\|w\|_{W^{2,2}(B(x,1/2))}$ are uniformly bounded ($\forall x\in\mathbb{R}^{4}$).
+This follows from the interior estimates of [17, Theorem 4.9] applied to $w$ and $\Delta w=g(v)$, since we have seen that the norms $\|g(v)\|_{L^{2}(B(x,1))}$ are uniformly bounded.
+Next, an application of [17, Theorem 4.11] to $v$ and $\Delta v=w$, shows that the norms $\|v\|_{W^{4,2}(B(x,1/4))}$ are uniformly bounded ($\forall x\in\mathbb{R}^{4}$), since the norms $\|v\|_{L^{2}(B(x,1))}$ are uniformly bounded.
+Therefore, we deduce from the Sobolev embeddings that $v$ as well as its first derivatives are bounded and uniformly continuous on $\mathbb{R}^{4}$.
+To prove that $v$ converges asymptotically to 0, we first establish that the set $\{|v|>\epsilon\}$ is bounded.
+Indeed, if $|v(x_{n})|>\epsilon$ holds for a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, then in view of the uniform continuity of $v$, there exists $\delta>0$ such that $|v|>\epsilon/2$ holds on the balls $B(x_{n},\delta)$, which by passing to a subsequence can be assumed to be disjoint.
+Consequently, we obtain that $\mu(\{|v|>\epsilon/2\})=\infty$, a contradiction.
+Now, suppose that there exists a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, and $|v(x_{n})|>\eta$, for some $\eta\in(0,\epsilon)$.
+According to what precedes, we have $|v(x_{n})|\in[\eta,\epsilon]$ (for $n$ large enough).
+Then, in view of the uniform continuity of the function $\mathbb{R}^{4}\ni x\mapsto G(v(x))$ and (5), there exists $\delta>0$, such that $|G(v)|>\beta>0$ holds on the balls $B(x_{n},\delta)$, for some $\beta>0$.
+As previously, the balls $B(x_{n},\delta)$ can be assumed to be disjoint, thus we get $G(v)\notin L^{1}(\mathbb{R}^{4})$, a contradiction.
+Therefore, we proved that $\lim_{|x|\to\infty}v(x)=0$.
+In a similar way, one can see that $\lim_{|x|\to\infty}|\nabla v(x)|=0$.
+(iii) To derive Pohozaev identity, reproduce the arguments in the proof of [23, Theorem 1.2]. ∎
 Finally, we establish that $D^{2}u$ is interchangeable with $\Delta u$ in the expression of functional $K$.
@@
 $=$ $\int_{\mathbb{R}^{4}}|\Delta\phi|^{2}+\lim_{r\to\infty}\sum_{i,j=1}^{4}\Big[\int_{\partial B_{r}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big],$
-where $B_{r}=\{x\in\mathbb{R}^{4}:|x|<1\}$, and $\nu=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})$ is the outward unit normal vector at $\partial B_{r}$.
+where $B_{r}=\{x\in\mathbb{R}^{4}:|x|<r\}$, and $\nu=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})$ is the outward unit normal vector at $\partial B_{r}$.
 Moreover, since $|\nabla\phi||D^{2}\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}=0$.
 Therefore, we conclude that
-$\Big|\int_{\partial B_{r_{k}}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
+$\Big|\int_{\partial B_{r_{k}}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r_{k}^{\frac{3}{4}})\to 0,$
 as $k\to\infty$, and $\int_{\mathbb{R}^{4}}|D^{2}\phi|^{2}=\int_{\mathbb{R}^{4}}|\Delta\phi|^{2}$.
@@
 On the one hand, it is clear that $\lim_{r\to\infty}\int_{B_{r}}\phi\cdot\Delta\phi=\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$.
-On the other hand, we notice that since $|\nabla\phi||\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi=0$.
+On the other hand, we notice that since $|\nabla\phi||\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}=0$.
 Therefore, we have
-$\Big|\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
+$\Big|\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r_{k}^{\frac{3}{4}})\to 0,$
 as $k\to\infty$, and $\lim_{k\to\infty}\int_{B_{r_{k}}}|\nabla\phi|^{2}=-\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$.
@@
 Consequently, it holds $\int_{\mathbb{R}^{4}}|\nabla\phi_{n}|^{2}=-\int_{\mathbb{R}^{4}}\phi_{n}\cdot\Delta\phi_{n}$, and as $n\to\infty$, we get that $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $\int_{\mathbb{R}^{4}}|\nabla u|^{2}=-\int_{\mathbb{R}^{4}}u\cdot\Delta u$.
+Remark 6.
+It follows from Lemma 11 that given the potential $G(u)=|u|^{2}\ln|u|$, we have $\mathcal{C}=\{u\in H^{2}(\mathbb{R}^{4}):\,G^{-}(u)\in L^{1}(\mathbb{R}^{4})\}$.
+Indeed, if $u\in\mathcal{C}$, it is clear that $u\in L^{2}(\mathbb{R}^{4})$, since $G(u)\in L^{1}(\mathbb{R}^{4})$, thus Lemma 11 implies that $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $u\in H^{2}(\mathbb{R}^{4})$.
+Conversely, if $u\in H^{2}(\mathbb{R}^{4})$, we have $u,\,\nabla u\in H^{1}(\mathbb{R}^{4})\subset L^{4}(\mathbb{R}^{4})$, and $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$, since $u\in L^{2}(\mathbb{R}^{4})$.
+Moreover, it holds $G^{+}(u)\in L^{1}(\mathbb{R}^{4})$, in view of the bound $G^{+}(u)\leq|u|^{4}$.
 5 Conclusion
@@
 $d\geq 2$ for (1), and $d\geq 4$ for (2)).
-In contrast with the second order system (1), none asymptotic condition on the potential $G$ is required for the fourth order system (2), since for the latter, the relevant class $\cal C$ of functions is contained in $L^{\infty}$.
+In dimension $d=2$, the ground state solution of the scalar second order equation (1) ($m=1$) was constructed in [7], assuming that the potential $G$ satisfies an exponential growth condition.
+This hypothesis is required to obtain compactness properties from the Trudinger inequality.
+Similarly, for the fourth order system (2) in dimension $d=4$, an exponential growth condition is assumed for $G$.
 We also point out that if the space dimension is smaller than that of the limiting case, then the problem of finding nontrivial solutions converging asympototically to 0 is different.